### v1

# Detecting and Filtering Unsafe Training Data via Data Attribution

Abstract
Large language models (LLMs) are vulnerable to unsafe training data that even small amounts of unsafe data can lead to harmful model behaviors. Detecting and filtering such unsafe training data is essential for trustworthy model development. Current state-of-the-art (SOTA) approaches typically rely on training moderation classifiers which requires significant computational overhead and are limited to predefined taxonomies, making them less adaptable to evolving safety concerns. Moreover, these classifiers lack insight into the training process, limiting their effectiveness in filtering unsafe data. To address these limitations, we propose DABUF, leveraging data attribution to detect and filter unsafe training data by attributing harmful model outputs to influential training data points. DABUF enables flexible identification of various unsafe data types without predefined taxonomies. However, in practice, model outputs can be complex with combined safe linguistic features and unsafe content, leading to reduced attribution accuracy. In such cases, DABUF will integrate moderation classifiers to identify a minimal subset of unsafe training data for targeted attribution (such as jailbreak). When model outputs are relatively straightforward, DABUF uses model outputs directly as the attribution targets. We evaluate the performance on two different tasks: in filtering jailbreaking training data and in identifying and mitigating gender bias. DABUF outperforms SOTA approaches by up to 7.5% in detection AUPRC in jailbreaking scenarios, and 44.1% in detecting gender bias. Moreover, retraining on DABUF-filtered data leads to higher model safety across experiments, underscoring its versatility in addressing a broad spectrum of unsafe data issues.
Detecting and Filtering Unsafe Training Data via Data Attribution
Yijun Pan1 Taiwei Shi2 Jieyu Zhao2 Jiaqi W. Ma3 1University of Michigan 2University of Southern California 3 University of Illinois Urbana-Champaign
1 Introduction
Large Language Models (LLMs) are known to exhibit various unsafe behaviors, including toxicity, stereotyping, privacy leaks, and ethical violations (Wang et al., 2024a). A primary source of these issues is unsafe training data (Jiang et al., 2024; Chen et al., 2024). For instance, inherent biases or toxic content in a dataset can lead to harmful responses (Jiang et al., 2024; Ouyang et al., 2022), while deliberate attacks, such as adversarial prompts or injected backdoors, can be used to bypass safety alignments (Chen et al., 2024; Zou et al., 2023; Li et al., 2024b). Consequently, identifying and removing these unsafe training instances is critical for mitigating risks and building safer LLMs.
Existing methods for detecting and filtering unsafe training data typically rely on moderation classifiers. Online API tools, such as the Perspective API  [^1] https://www.perspectiveapi.com/ and OpenAI’s Moderation API (Markov et al., 2023), focus on certain predefined toxicity taxonomies, but struggle to generalize to nuanced and emerging unsafe artifacts beyond these predefined taxonomies Weber et al. (2025). Fine-tuned detection models, including Llama-Guard-3-8B (Llama Team, 2024) and Wildguard (Han et al., 2024), require significant time and resources for additional data collection and training. Moreover, these moderation classifiers primarily detect semantically unsafe training data without considering the influence of each data point on model training, resulting in suboptimal filtering effectiveness for enhancing model safety.
In this work, we introduce Data-Attribution-Based Unsafe Training Data Detection and Filtering (DABUF), a method that leverages data attribution techniques to enhance unsafe data detection and filtering in LLMs. Data attribution is a family of methods that quantify the influence of individual training data points on specific model outputs (Koh and Liang, 2020; Pruthi et al., 2020). Our central hypothesis is that unsafe training data exerts greater influence on unsafe outputs; thus, attributing these outputs back to their influential training instances can reveal which data points are responsible. These methods do not require additional training data and can be applied flexibly to diverse types of unsafe model outputs.
However, applying data attribution naively—by directly attributing unsafe model generations to their training data—is ineffective in many cases. LLM outputs, particularly in long-form generations such as jailbreaking attacks, are influenced by a mix of benign and unsafe training data. Since model generations include common linguistic structures (e.g., stop words, neutral phrases) alongside unsafe content, direct attribution to the entire sequence leads to a noisy attribution signal, reducing the precision of unsafe data identification.
To address this challenge, our method introduces a targeted filtering mechanism. Specifically, for long-form outputs, such as those found in jailbreaking scenarios, we first use moderation classifiers to identify a small subset of clearly unsafe training data. We then use this subset as the attribution target, allowing us to refine the attribution process and isolate the most influential unsafe training instances. In contrast, for shorter model outputs, such as those in gender bias scenarios, where the influence of training data is more direct and less noisy, standard data attribution techniques are sufficient without additional moderation filtering.
We validate our approach through experiments in two distinct setups. In jailbreaking scenarios involving adversarial prompts that lead to noisy or long unsafe outputs, we apply the proposed DABUF with moderation classifiers for initial unsafe data identification. Conversely, in the gender bias scenario—where outputs are relatively concise—we directly apply DABUF on the model generation. Experimental results show that our approach achieves superior detection performance across different model architectures in jailbreaking scenarios and deliver improved safety when models are retrained with filtered data, outperforming state-of-the-art detection methods. Furthermore, our method generalizes effectively to gender bias scenarios, highlighting its versatility.
2 Related work
2.1 Sources of Unsafe Training Data in LLMs
Recent studies (Yi et al., 2024; Qi et al., 2023) reveal that malicious fine-tuning can severely compromise safety alignment, even with limited exposure to unsafe data. Unfortunately, current online fine-tuning services are inadequate at detecting these unsafe training data, leaving LLMs vulnerable to potential exploitation (Qi et al., 2023).
Unsafe data may also emerge from synthetic training data generation. For instance, Wang et al. (2022) generate samples by conditioning LLMs on specific keywords and target labels, while Wang et al. (2023) synthesize fine-tuning data from LLM-generated responses. However, as recent safety research (Wang et al., 2024a) indicates, even highly aligned models like GPT-4 and GPT-3.5 exhibit unsafe behaviors, suggesting that synthetic data can introduce significant risks.
In addition, inherent biases in training data pose challenges that current detection methods are not equipped to handle. Studies have shown that gender bias in training data can lead LLMs to develop skewed assumptions about occupations (Kotek et al., 2023), while cognitive biases during training undermine model reliability in high-stakes decisions (Itzhak et al., 2024; Echterhoff et al., 2024).
The diverse source of unsafe training data highlights the need for more flexible and adaptable detection methods. Current moderation classifiers, often designed for specific content moderation tasks, are insufficient for addressing the complexity and variability of unsafe data in training pipelines.
2.2 Unsafe Training Data Detection in LLMs
Existing efforts to detect unsafe training data primarily focus on content moderation classifiers. For example, online moderation tools such as OpenAI’s Moderation API (Markov et al., 2023) are developed to detect harmful content. Recently, there has been growing efforts in developing LLM-based classifiers. One line of research has explored fine-tuning open-source LLMs on specifically curated safety dataset to develop moderation classifiers. Examples of such classifiers include Llama-Guard-3-8B (Llama Team, 2024), Wildguard (Han et al., 2024), Aegis-Guard (Ghosh et al., 2024), and ShieldGemma (Zeng et al., 2024). Another line of research focuses on leveraging LLMs directly as judges for unsafe data detection (Franco et al., 2023; Li et al., 2024a). For instance, SafetyAnalyst (Li et al., 2024a) proposes using LLMs to generate “harm-benefit” tree for interpretable content moderation.
Beyond content moderation classifiers, some recent studies have leveraged the internal structures of models for unsafe data detection. For example, GradSafe (Xie et al., 2024) utilizes gradient similarity with respect to safety-critical parameters to identify unsafe data, while BEBC (Zheng et al., 2024) employs fine-tuned BERT embeddings for content moderation.
Unlike online moderation tools and LLM-based classifiers, our methods eliminate the need for data curation and additional training. Furthermore, by operating independently of predefined safety taxonomies, our approach demonstrates greater flexibility and can effectively handle a wider range of safety-related scenarios. In contrast to other non-classifier approaches, such as GradSafe, our methods address the problem through the lens of data attribution. This perspective enhances detection performance by incorporating unsafe model behavior and capturing the relationship between training data and model outputs, ultimately fostering safer model developments.
2.3 Data Attribution for LLMs
Data attribution methods aim to quantify the impact of individual training samples on a model’s predictions for specific test cases (Koh and Liang, 2020). These methods have the potential to detect unsafe training data, as such data are likely to exert a disproportionate influence on unsafe model outputs, making them distinguishable from the broader benign dataset. Recently a variety of data attribution methods have been proposed to estimate the influence of training data in the context of LLMs. These include gradient-based methods (Xia et al., 2024; Kwon et al., 2024), simulator-based methods (Guu et al., 2023; Chai et al., 2024) and game-theoretic methods (Wang et al., 2024b, c). Estimated influence scores have been utilized for tasks such as identifying mislabeled data (Pruthi et al., 2020), understanding memorization (Feldman and Zhang, 2020), and data valuation (Choe et al., 2024). While data attribution methods have various applications in LLMs, they are computationally intensive, limiting their applicability to larger models. Despite recent advancements in efficient influence estimation methods Kwon et al. (2024), the computational burden remains a challenge. Gradient-similarity-based approaches, as highlighted in previous works Xia et al. (2024); Pruthi et al. (2020), offers an efficient solution, making it better suited for scaling to LLMs.
3 Data-Attribution-Based Unsafe Training Data Detection and Filtering
Our proposed method consists of two phases: detection and filtering. The primary technical challenge arises in the detection phase, where we identify unsafe data points in the training dataset that contribute to unsafe model behaviors. In the filtering phase, we mitigate these behaviors by removing the data points most likely to be unsafe.
3.1 Unsafe Training Data Detection
We first describe the problem setup of unsafe training data detection. Consider a training dataset with a mixture of benign and unsafe data:
$\mathcal{D}_{\text{train}}=\mathcal{D}_{\text{benign}}\cup\mathcal{D}_{\text{unsafe}},$
where $\mathcal{D}_{\text{benign}}$ refers to the benign dataset that is safe to train on while $\mathcal{D}_{\text{unsafe}}$ is the unsafe training dataset that could lead to unsafe model behaviors. In addition, we assume access to a (small) target dataset $\mathcal{D}_{\text{target}}$ that consists of unsafe model outputs or examples of the unsafe training data. In practice, these examples may come from user reports or manual inspection of a small portion of training data.
The goal of unsafe training data detection is to retrieve $\mathcal{D}_{\text{unsafe}}$ from the entire $\mathcal{D}_{\text{train}}$ with high precision and recall, possibly using the information from the target set $\mathcal{D}_{\text{target}}$. A high-quality detection method will help us obtain a cleaner training dataset without overly removing safe training data in the filtering phase.
Data-Attribution-Based Detection
We propose to detect the unsafe training data by measuring the influence of each training data point $z\in\mathcal{D}_{\text{train}}$ on the likelihood of the model generating the examples in the target dataset $\mathcal{D}_{\text{target}}$. Intuitively, since $\mathcal{D}_{\text{target}}$ consists of unsafe examples, a training data point with higher influence is more likely to be unsafe. Formally, we denote the influences as
$\text{Inf}(z,\mathcal{D}_{\text{target}}),\quad z\in\mathcal{D}_{\text{train}}.$
In this work we use gradient similarity (Pruthi et al., 2020) to efficiently estimate training data’s influence on model generations, which is a scalable method that has been widely used in data attribution for LLMs (Xia et al., 2024). Consider a model parameterized by $\theta$, and denote the negative log-likelihood as $\ell(\cdot;\theta)$. The influence of a training data point $z\in\mathcal{D}_{\text{train}}$ on the target dataset $\mathcal{D}_{\text{target}}$ is defined as following:
$\mathrm{Inf}(z,\mathcal{D}_{\text{target}}):=\eta\cos(\nabla\ell(\mathcal{D}_{\text{target}};\theta),\nabla\ell(z;\theta)),$
where $\eta$ is the average learning rate, $\nabla\ell(\mathcal{D}_{\text{target}};\theta)$ is the gradient of the negative log-likelihood with respect to $\theta$ evaluated on the target set $\mathcal{D}_{\text{target}}$, and $\nabla\ell(z;\theta)$ is the gradient evaluated on the training point $z$. To improve computational efficiency in practice, we follow Xia et al. (2024) to reduce the gradient dimension to $d=8192$ via random projection, and adopt optimizer-aware training gradients.
Data Attribution for LLM Outputs
In the context of LLMs, the model output is a sequence of tokens. Specifically, for each example $x\in\mathcal{D}_{\text{target}}$, it can be represented as $x=(p,r)$, where $p$ is the input prompt and $r$ is the output response, both of which are a sequence of tokens. Existing literature (Xia et al., 2024) typically defines $\nabla\ell(\mathcal{D}_{\text{target}};\theta)$ as
$\nabla\ell(\mathcal{D}_{\text{target}};\theta)=\sum_{x\in\mathcal{D}_{\text{target}}}\nabla\log p(r|p;\theta),$
where one can further expand the conditional probability $p(r|p;\theta)$ defined by an autoregressive LLM as following:
$\nabla\ell(r;\theta,p)=\sum_{i=1}^{|r|}\nabla\ell(r_{i}|p,r_{<i};\theta).$
However, we find that naively applying the data attribution method defined above to long-form unsafe outputs often yields suboptimal performance for identifying unsafe training data. When a response $r$ contains many benign tokens and only a few segments of genuinely unsafe content, the attribution signal becomes diluted by the larger volume of neutral or benign tokens. In particular, the tokens directly associated with unsafe content–henceforth referred to as unsafe tokens–carry disproportionately stronger gradients, indicating their direct link to harmful outputs. Yet, when the model’s overall attribution signal is aggregated over all tokens, these critical unsafe signals become overwhelmed by the contributions of benign tokens, resulting in noisy and less precise detection of unsafe training data points. This imbalance is especially problematic in long-form scenarios like jailbreaking attacks, where substantial portions of the model response may be benign filler text interspersed with targeted unsafe content. Consequently, the naive approach of attributing every token in the entire generation fails to isolate the truly influential unsafe training instances. Please refer to Appendix B.1 for more detailed empirical evidence of this observation.
Leveraging Externally-Identified Unsafe Data for Effective Attribution
To address the aforementioned issue in scenarios where the model outputs are long, we propose to take ground-truth labels (instead of the model outputs) from a small, externally identified subset of unsafe training data as the target dataset to attribute. For example, in the jailbreaking scenario, we first an LLM-based classifier (Llama-3-Guard-8B (Llama Team, 2024)) to screen the entire training dataset and obtain a small candidate set of unsafe training data $\mathcal{D}_{\text{cand}}$. Because such classifiers can have high false-positive rates, we further perform human annotation on $\mathcal{D}_{\text{cand}}$ to filter out benign data points, resulting in a smaller, verified unsafe subset $\mathcal{D}_{\text{identified}}$. The target set is then set as $\mathcal{D}_{\text{target}}=\mathcal{D}_{\text{identified}}$. In all of our experiments, $|\mathcal{D}_{\text{cand}}|$ is well below 200–manageable for human inspection–whereas the full training set exceeds 40,000 samples.
3.2 Unsafe Training Data Filtering
Using the estimated influence scores, we perform retrieval to identify the training examples most influential on the unsafe target samples. We select the top $K$ elements from the ranked list of $\text{Inf}(z,\mathcal{D}_{\text{target}})$ values. Let $\mathcal{S}_{K}(\mathcal{D}_{\text{target}})\subseteq\mathcal{D}_{\text{train}}$ denote the set of $K$ most influential training samples selected, when the target dataset $\mathcal{D}_{\text{target}}$ is used as the target for attribution. By removing $\mathcal{S}_{K}(\mathcal{D}_{\text{target}})$ from the training data, we construct a cleaner dataset that is expected to improve model safety upon retraining.
4 Experiments: Jailbreaking Injection Detection
In this section, we evaluate the proposed method in the jailbreaking injection detection scenario. Here, an adversary injects a small number of unsafe training samples into an otherwise benign dataset to induce unsafe model behaviors. Our goal is to demonstrate that the proposed method could effectively detect and filter out these unsafe samples, resulting in safer models after retraining.
4.1 Experimental Setup
Overview.
We focus on a realistic training scenario wherein the benign portion of the data can be heterogeneous, consisting of:
  1. 1.
    Fully benign prompt-response pairs with no harmful content.
  2. 2.
    Safe demonstrations: pairs where the prompt may be unsafe, but the response is aligned and refuses or mitigates the request. These “safe” demonstrations have been shown to improve model safety (Jain et al., 2023). We denote the set of safe demonstrations as $\mathcal{D}_{\text{safe}}$.
We augment this benign dataset with a small set of unsafe injections designed to induce harmful responses. Across all experiments, the total injection ratio is below $0.025\%$. Identifying these few harmful training instances is extremely challenging, but crucial for mitigating jailbreaking vulnerabilities.
Datasets
We use the dataset Ultrachat 200k as our benign dataset and consider two unsafe datasets, ToxicChat and XSTest-Response. For each unsafe dataset, we split the dataset into $\mathcal{D}_{\text{safe}}$ (unsafe prompts with safe responses), $\mathcal{D}_{\text{unsafe}}$ (unsafe prompts and unsafe responses), and $\mathcal{D}_{\text{test}}$ (held-out unsafe prompts for evaluation). In our experiments, we inject both the $\mathcal{D}_{\text{safe}}$ and $\mathcal{D}_{\text{unsafe}}$ from the unsafe dataset into the benign Ultrachat 200k dataset to form the whole training dataset $\mathcal{D}_{\text{train}}=\mathcal{D}_{\text{benign}}\cup\mathcal{D}_{\text{safe}}\cup\mathcal{D}_{\text{unsafe}}$.
  • Ultrachat 200k. Ultrachat 200k [^2] https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k is a heavily filtered version of the UltraChat Ding et al. (2023) dataset, which comprises over 200k instructions for instruction fine-tuning purposes. We use a subset of the train sft split for our $\mathcal{D}_{\text{benign}}$, comprising $41,573$ samples.
  • ToxicChat. ToxicChat [^3] https://huggingface.co/datasets/lmsys/toxic-chat Lin et al. (2023) is a dataset consisting prompt-response pairs annotated with prompt toxicity and jailbreak-ness (response toxicity), curated from user interactions. We use a subset of the latest version of ToxicChat: ToxicChat-0124, and apply the following split: $|\mathcal{D}_{\text{safe}}|=127$, $|\mathcal{D}_{\text{unsafe}}|=97$, and $|\mathcal{D}_{\text{test}}|=30$.
  • XSTest-Response. XSTest-Response [^4] https://huggingface.co/datasets/allenai/xstest-response Han et al. (2024) is a dataset consisting of responses to the original XSTest Röttger et al. (2023), which includes 446 annotations of prompt harmfulness and model response harmfulness. We use the official response_harmfulness subset and apply the following split: $|\mathcal{D}_{\text{safe}}|=121$, $|\mathcal{D}_{\text{unsafe}}|=65$, and $|\mathcal{D}_{\text{test}}|=20$.
4.2 Evaluation Metrics
Given the retrieved set $\mathcal{S}_{K}(\mathcal{D}_{\text{target}})$ containing $K$ top influential training data to the validation set, we define the precision and recall as:
precision $=\frac{|\mathcal{S}_{K}(\mathcal{D}_{\text{target}})\cap\mathcal{D}_{\text{unsafe}}|}{|\mathcal{S}_{K}(\mathcal{D}_{\text{target}})|}$
recall $=\frac{|\mathcal{S}_{K}(\mathcal{D}_{\text{target}})\cap\mathcal{D}_{\text{unsafe}}|}{|\mathcal{D}_{\text{unsafe}}|}$
We adopt the Area Under Precision Recall Curve (AUPRC) as well as the precision, recall and F1 scores calculated with the top 100 samples identified for a comprehensive evaluation of baselines models and our methods.
To evaluate model safety, we employ Attack Success Rate (ASR) on the test set $\mathcal{D}_{\text{test}}$, which measures the proportion of the unsafe prompts in $\mathcal{D}_{\text{test}}$ that successfully elicit unsafe responses from the model.
4.3 Baselines
We include baselines from three categories: online API tools (OpenAI moderation API), fine-tuned LLM as detectors (Llama-Guard-3-8B, Wildguard) and other model-free methods (GradSafe).
OpenAI moderation.
The OpenAI Moderation API Markov et al. (2023) is an online moderation tool that assess whether the content is unsafe across 11 safety genres. We take the binary prediction label from the model to calculate precision, recall and F1. For AUPRC we use the model’s highest confidence across all safety genres.
Llama-Guard-3-8B.
Llama-Guard-3-8B Llama Team (2024) is a Llama-3.1-8B pretrained model, fine-tuned for content safety classification. For AUPRC we use the probability of outputting the token “unsafe”, consistent with previous methodologies Xie et al. (2024).
Wildguard.
Wildguard Han et al. (2024) is an open one-stop moderation model trained on a Mistral-7B model that detects prompt harmfulness, response harmfulness, and whether the response is a refusal to the prompt. Similar to Llama-Guard-3-8B, we use the probability of outputting the token “unsafe” as the confidence to calculate AUPRC.
GradSafe.
GradSafe Xie et al. (2024) differs fundamentally in methodology by analyzing gradients with respect to safety-critical parameters of Llama-2, specifically focusing on the gradient of the model’s compliance response to prompts. In contrast, our approach directly traces unsafe behaviors back to the training data by leveraging token-level attributions to identify the sources of unsafe outputs. GradSafe operates independently of the model’s responses, providing pre-hoc moderation akin to LLM classifiers, whereas our method emphasizes post-hoc attribution to uncover the origins of unsafe model behavior.
4.4 Results and Discussion
In this section, we present the results of baseline methods and our approach for detecting and filtering jailbreaking data.
We first demonstrate that fine-tuning language models on jailbreaking training data can effectively compromise their safety. Table 1 presents the ASR across different models and datasets. In comparison to training on the benign dataset $\mathcal{D}_{\text{benign}}$ only, the ASR is significantly higher when training with $\mathcal{D}_{\text{train}}$ that consists of unsafe training data, confirming that training on injected unsafe data results in unsafe model behaviors.
Table 1: Attack Success Rate (ASR) across trained models and datasets. A higher ASR indicates a more unsafe model.
Model Data ToxicChat XSTest-Response
Llama-3-8B $\mathcal{D}_{\text{train}}$ 93.3% 50%
$\mathcal{D}_{\text{benign}}$ 66.7% 0%
Gemma-2-9B $\mathcal{D}_{\text{train}}$ 90.0% 100%
$\mathcal{D}_{\text{benign}}$ 83.3% 70%
Table 2: AUPRC of baseline models and our method. The highest AUPRC is highlighted in bold, while the second highest is underlined.
Method ToxicChat (%) XSTest-Response (%)
OpenAI Moderation API 11.7 11.8
Llama-Guard-3-8B 30.3 82.5
Wildguard 44.5 85.9
GradSafe 30.7 47.3
Llama-3-8B-DABUF 52.0 74.1
Gemma-2-9B-DABUF 49.1 64.1
Table 3: Precision, recall, and F1 scores of baseline models and our method, calculated based on the top 100 identified training data points. The highest F1 score is highlighted in bold, and the second highest is underlined.
ToxicChat XSTest-Response
Method Precision (%) Recall (%) F1 (%) Precision (%) Recall (%) F1 (%)
OpenAI Moderation API 16.0 16.5 16.2 19.0 28.8 22.9
Llama-Guard-3-8B 33.0 34.0 33.5 57.0 86.4 68.7
Wildguard 46.0 47.4 46.7 59.0 89.4 71.1
GradSafe 51.1 47.4 49.2 45.0 68.2 54.2
Llama-3-8B-DABUF 51.0 52.6 51.8 57.0 86.4 68.7
Gemma-2-9B-DABUF 52.0 53.6 52.8 53.0 80.3 63.9
Table 4: Attack Success Rate (ASR) comparison between models retrained with the top 100 unsafe training samples filtered by baseline methods and DABUF. A higher ASR reflects a more unsafe model.
Model Filtering Method ToxicChat XSTest-Response
Llama-3-8B OpenAI Moderation 96.7% 25%
GradSafe 93.3% 25%
Wildguard 90.0% 15%
Llama-Guard-3-8B 90.0% 50%
DABUF 86.7% 5%
Gemma-2-9B OpenAI Moderation 90% 45%
GradSafe 86.7% 60%
Wildguard 80% 20%
Llama-Guard-3-8B 90% 30%
DABUF 66.7% 15%
Table 2 and 3 respectively show the AUPRC and the top 100 precision, recall, and F1 of unsafe training data detection across models and methods. Our method demonstrates superior or comparable performance across different experimental settings. Notably, DABUF applied to Llama-3-8B achieves the highest performance in the ToxicChat injection experiment, showcasing the advantages of using data attribution for detecting unsafe training data. While state-of-the-art classifiers like Wildguard outperform DABUF in XSTest-Response experiments, this is largely due to XSTest-Response’s focus on toxicity and explicit harm, aligning with their in-distribution training data.
More importantly, Table 4 highlights that the proposed DABUF significantly outperforms all baseline methods in terms of the ASR of models retrained after filtering out the top 100 unsafe training samples identified by different methods. The results demonstrate that our data attribution approach, which explicitly accounts for the model training process, effectively identifies and filters unsafe training data that contributes the most to the unsafe model behaviors of interest, which leads to models with better safety when retrained on filtered datasets.
5 Experiments: Gender Bias Mitigation
In this section, we further evaluate the proposed method in a gender bias mitigation scenario, where most moderation classifiers are not applicable.
5.1 Problem Setup
Prior research (An et al., 2024) has shown that LLMs can exhibit gender biases, particularly in contexts such as hiring decisions. To explore this issue, we present a scenario where training data contains inherent gender biases. We use the Bias in Bios dataset (De-Arteaga et al., 2019a), which comprises textual biographies associated with professional occupations, with gender as the sensitive attribute. From the training split, we sampled a subset of $10,000$ biography-occupation pairs ($\mathcal{D}_{\text{benign}}$) and injected it with 150 biased biography-occupation pairs ($\mathcal{D}_{\text{unsafe}}$) to form the training set ($\mathcal{D}_{\text{train}}$). For the target set, we generated gender-biased model responses for 50 occupation prediction prompts ($\mathcal{D}_{\text{target}}$). Additional details of the experiment are provided in Appendix A. Since model responses in this scenario are relatively simple, we do not use externally identified unsafe training data as we did in jailbreaking setups.
5.2 Evaluation Metrics
Similar with jailbreaking injection, we evaluate the detection performance via precision, recall, F1 score, and AUPRC.
To evaluate gender bias in trained models, we adopt the metric of True Positive Rate (TPR) Gender Gap following De-Arteaga et al. (2019b). Let $A\in\{0,1\}$ be the gender attribute, where $A=0$ represents male and $A=1$ represents female. Additionally, $Y\in\{0,1\}$ denotes the occupation, with $Y=1$ indicating the person is a physician and $Y=0$ indicating the person is a nurse. In a held-out test set where the ground truth labels are all physicians, the TPR Gender Gap is defined as:
$\text{Gap}=TPR_{A=0}-TPR_{A=1},$
where $TPR_{A=a}=p(\hat{Y}=1|A=a,Y=1)$ is the TPR for the gender group $A=a$. The TPR Gender Gap intuitively quantifies the extent to which a model’s predictions favor physician against nurse when the gender-related information in the prompt is switched from female to male. A larger TPR Gender Gap observed on the held-out test set indicates that the model exhibits stronger gender biases.
5.3 Results and Discussion
First, we demonstrate that training a model on gender-biased data leads to behaviors that reflect gender biases. As can be seen in Table 5, models trained on gender-biased data exhibit a significant higher TPR Gender Gap in comparison to those trained on unbiased data.
Table 5: TPR Gender Gap on unbiased and biased datasets.
Model Dataset TPR Gender Gap
Llama-3-8B Unbiased 0.04
Biased 0.16
Gemma-2-9B Unbiased 0.02
Biased 0.08
Table 6: AUPRC and precision, recall, and F1 of the baseline model and our methods. The highest AUPRC and F1 value are highlighted in bold, while the second highest underlined. The suffix DABUF denotes our method.
Method AUPRC Precision Recall F1
Ada3 0.089 0.500 0.330 0.400
Llama-3-8B-DABUF 0.474 0.610 0.407 0.488
Gemma-2-9B-DABUF 0.530 0.670 0.447 0.536
Table 7: TPR Gender Gap comparison between models retrained with the top 100 biased samples filtered by Ada3 and our method. A higher TPR Gender Gap indicates a model with more gender bias. The lowest value is highlighted in bold.
Model Filtering Method TPR Gender Gap
Llama-3-8B None 0.16
Ada3 0.28
DABUF 0.14
Gemma-2-9B None 0.08
Ada3 0.04
DABUF 0.02
Table 6 presents the AUPRC, precision, recall, and F1 score results. Embedding-based tools like Ada3 exhibit poor performance in detecting unsafe data, suggesting that general-purpose embedding tools are not well-suited for detecting gender-biased data. In contrast, as a model-free approach, our method adapts effectively to different injection scenarios using diverse unsafe validation data.
Table 7 presents the TPR Gender Gap for models retrained after removing the top 100 influential samples identified. These results demonstrate that our approach effectively detects training samples that contribute to gender bias, and retraining on the filtered data leads to models with reduced gender bias. Notably, while embedding-based methods like Ada3 achieve non-trivial detection performance, filtering using Ada3 actually results in a higher TPR Gender Gap compared to training on the original data for Llama-3-8B model. This suggests that embedding-based methods may not be suitable for training data filtering. We believe this occurs because embedding-based models focus primarily on semantics, causing them to remove both unsafe and safe training data indiscriminately. This indiscriminate removal undermines the model’s ability to effectively learn the nuances of the occupation prediction task.
6 Conclusion
This work proposes the challenge of detecting and filtering unsafe training data embedded within larger benign datasets. We propose using Data-Attribution-Based Unsafe Training Data Detection and Filtering (DABUF) to detect and filter unsafe training data that contributes to unsafe model behavior. In scenarios where model outputs are long, DABUF overcomes the noisy aggregation of token-wise gradients by utilizing externally identified unsafe training data for effective attribution. Remarkably, DABUF’s performance for jailbreaking samples detection and filtering surpasses those by SOTA LLM-based moderation classifiers across various models and datasets. Furthermore, our method is not confined to existing unsafe taxonomies, demonstrating its adaptability to broader unsafe scenarios, such as gender bias mitigation.
Limitation
In this work, we proposed an effective and versatile method to detect unsafe training data in a realistic scenario. However, we acknowledge several avenues for future improvements.
Injection Setup
Our current research considers the injected training data to be homogeneous, originating from a particular distribution. However, real-world scenarios likely involve more heterogeneous injected training data, with each data point potentially exerting different influences on various genres of unsafe model behavior. This complexity suggests the need for more nuanced injection strategies in future research.
Detection Taxonomy
Our current work does not provide a fine-grained detection taxonomy, as both training and validation data encompass multiple unsafe genres. While we recognize that carefully selecting validation data could potentially enable detection of specific genres, we consider this beyond the scope of the present study and recommend it as a promising direction for future investigation.
Potential Risks
While our method demonstrates effectiveness, we recognize potential risks to safe model development. Recent research on adversarial attacks in data attribution Wang et al. (2024d) suggests significant vulnerabilities in influence estimation techniques. Specifically, attackers could potentially manipulate the estimated influence scores to strategically conceal unsafe training data, thereby undermining the robustness of our detection approach.
Appendix A Gender Bias setup
In this section we illustrate the construction of gender bias injection from the Bias in bios De-Arteaga et al. (2019b) dataset.
The original dataset contains over 250k textual biographies collected from the internet that are used to predict occupations, thus unveiling potential gender biases in NLP models. An example biography is as follows:
Example data from Bias in bios Biography:
She is able to assess, diagnose and treat minor illness conditions and exacerbations of some long term conditions. Her qualifications include Registered General Nurse, Bachelor of Nursing, Diploma in Health Science, Emergency Care Practitioner and Independent Nurse Prescribing.
Occupation:
Nurse
Therefore we use the following prompt format to fine-tune language models on occupation prediction:
Prompt User: Given the following biography of a person, predict the occupation that person has: {biography} Assistant: {occupation}
Note that in the biography there exists gender indicators that hint about the person’s gender, example of gender indicators include: he/she, him/her, his/her, Ms/Mr., etc.. We use the same gender indicator transform strategy of the original paper De-Arteaga et al. (2019b).
Following the discussion in De-Arteaga et al. (2019b), we formulate potential gender bias of language models to be:
Model’s tendency to associate certain professions to certain gender indicators with the same biography.
Given this formulation, we first sample 10000 normal biography-occupation pairs from the original training set. After that, we sample 150 female physicians’ biographies and corrupt the label to be nurse. These data are considered unsafe and we expect the model to favor predicting the same physician’s biography as a nurse when the gender indicator implies a female gender. Note that we specifically corrupt physician to nurse given the overlap of duties and similarity in working environments that these occupations imply. An example of injected data is:
Example of injected gender bias data User: Given the following biography of a person, predict the occupation that person has: Dr. Ho attended the University of Pennsylvania School of Medicine. Dr. Ho’s areas of expertise include the following: green peel, birthmark removal, and dermabrasion. Patients gave her an average rating of 2.0 stars out of 5. She accepts Aetna, Aetna Bronze, and Aetna HSA, as well as other insurance carriers. She is professionally affiliated with Jeanes Hospital. Assistant: physician $\rightarrow$ nurse
Note that in the original 10000 training set there are also normal entries for female/male nurses and physicians, so we consider the retrieval of these injected data as non-trivial.
A.1 Baselines
Since existing LLM safety classifiers are not adapted to detect gender biases, we employ the following models baselines.
Embedding Models
Because unsafe training data may be semantically distinct from benign training data, a detection approach that compares their representations provides a natural baseline. Specifically, we consider Ada3  [^5] https://openai.com/index/new-and-improved-embedding-model/, a SOTA model for texual embeddings. We compute embedding similarities between individual training examples and the mean embedding of the validation set, then identify training samples that exhibit the highest similarity scores.
Appendix B Analysis
B.1 Noisy Token-wise Gradients
In this section we show that model outputs can be noisy and thus directly using responses’ gradient results in sub-optimal results. A key observation is that training data impact the predictions of different tokens unevenly, with tokens containing unsafe content being the most affected. Therefore, unsafe tokens’ gradients result in greater influence when used for attribution compared to benign tokens’ gradients, making the overall gradient, as a summation of token-wise gradients, inherently noisy.
To elaborate this, we evaluate each token’s performance when used as target for attribution and show that token-wise gradients from model outputs have diverse performances for detection.
Specifically, we consider each token $t$’s detection performance by its detection precision:
$\text{prec}_{t}$ $=\frac{|\mathbf{S}_{M}(t)\cap\mathcal{D}_{\text{unsafe}}|}{|\mathbf{S}_{M}(t)|}$
By selecting the top $N$ token-wise gradients $t_{1},t_{2},\dots,t_{N}$ that achieve highest precision scores from the entire validation set, we used the filtered validation gradient as the target for attribution:
¯$\bar{\nabla}\ell(\mathcal{D}_{\text{val}};\theta)=\frac{1}{N}\sum_{i=1}^{N}\nabla\ell(t_{i};\theta)$
Figure 1: Performance analysis on ToxicChat of varying number of top-contributing tokens.
Figure 1 illustrates the impact of the number of included token-wise gradients on retrieval performance. Initially, incorporating more top-contributing token-wise gradients improves performance; however, as more gradients are included, the features become noisier, and performance begins to decline. This highlights the inherent noise in gradient features.
Table 8: Contribution of token’s gradient to detection for Llama-3-8B trained model in the XSTest-Response injection setting.
Token AUPRC
and 0.040
the 0.070
on 0.050
poison 0.176
commit 0.274
weapon 0.133
Table 8 shows the efficacy of individual tokens in identifying unsafe training data. As shown, common words such as and, the, and on are largely ineffective for detection, whereas explicit unsafe tokens such as poison, commit, and weapon make significant contributions. This disparity can be attributed to the mechanism behind unsafe training: unsafe training data has significant influence on the prediction of unsafe tokens while exerting minimal impact on common syntax tokens like and, the and on. Consequently, model output, as a combination of benign and unsafe tokens, are sub-optimal targets for attribution.
Table 9: AUPRC comparison between directly using model responses as the attribution target and DABUF.
Method ToxicChat (%) XSTest-Response (%)
Llama-3-8B-Response 4.70 35.4
Llama-3-8B-DABUF 52.0 74.1
Gemma-2-9B-Response 4.12 11.9
Gemma-2-9B-DABUF 49.1 64.1
Table 10: Precision, recall, and F1 scores comparison between directly using model responses as the attribution target and DABUF, calculated based on the top 100 identified training data points.
ToxicChat XSTest-Response
Method Precision (%) Recall (%) F1 (%) Precision (%) Recall (%) F1 (%)
Llama-3-8B-Response 14.0 14.4 14.2 19.0 28.8 22.9
Llama-3-8B-DABUF 51.0 52.6 51.8 57.0 86.4 68.7
Gemma-2-9B-Response 14.0 14.4 14.2 21.0 31.8 25.3
Gemma-2-9B-DABUF 52.0 53.6 52.8 53.0 80.3 63.9
This is further demonstrated by experimental results in Table 9 and table 10. DABUF, which leverages externally classified training data as its attribution target, consistently outperforms existing data attribution methods that rely solely on model responses. This finding suggests that depending exclusively on model responses can introduce noise, thereby hindering effective detection.
B.2 Validation Set Variety
In this section we delve into the effect of validation set variety on detection performance. Intuitively, having a larger validation set where model unsafe behaviors are observed across different domains helps with retrieval, especially when the training data contains multiple unsafe genres. By default we use a validation set of size 30 for retrieval in ToxicChat injection, where the injected data are in-the-wild user interactions with LLMs and thus contain diverse unsafe behaviors. By varying the size of $\mathcal{D}_{\textbf{identified}}$, we evaluate their retrieval performance.
Figure 2: Performance on ToxicChat with varying validation set size.
Figure 2 demonstrates the effect of validation set variety on unsafe training data detection. The sampling process for each size is repeated for 10 times and the mean as well as standard deviation are calculated accordingly. It is observed that a larger validation set size leads to higher detection performance. By including a more diverse validation set, we can capture a wider range of unsafe training data. Nevertheless, we note that by including as little as 20 validation samples, our approach already exceeds SOTA models like Llama-Guard-3-8B and Wildguard.
Appendix C Experimental Details
For experiments involving model training, we train for 4 epochs with a warm up ratio of 0.1 and used learning rate of 1e-4 for Llama-3-8B and 1e-5 for Gemma-2-9b respectively. The batch size is set to be 1 with no gradient accumulation. The training took place on Nvidia A40 GPUs. For training we use the official implementation of TRL von Werra et al. (2020) while for the calculation of AUPRC we use the official implementation from scikit learn [^6] https://scikit-learn.org/stable/. LoRA Hu et al. (2021) is used to reduce trainable parameters and decrease the size of the gradient features to accelerate gradient feature computation. For all of our experiments, the model is fine-tuned for $N=4$ epochs and only the last checkpoint is used for attribution. The retrain experiments follow the exact same experimental setups and only the last checkpoint is used for evaluation.

### v2

@@
-Detecting and Filtering Unsafe Training Data via Data Attribution
+Detecting and Filtering Unsafe Training Data via Data Attribution with Denoised Representation
 Abstract
-Large language models (LLMs) are vulnerable to unsafe training data that even small amounts of unsafe data can lead to harmful model behaviors.
-Detecting and filtering such unsafe training data is essential for trustworthy model development.
-Current state-of-the-art (SOTA) approaches typically rely on training moderation classifiers which requires significant computational overhead and are limited to predefined taxonomies, making them less adaptable to evolving safety concerns.
-Moreover, these classifiers lack insight into the training process, limiting their effectiveness in filtering unsafe data.
-To address these limitations, we propose DABUF, leveraging data attribution to detect and filter unsafe training data by attributing harmful model outputs to influential training data points.
-DABUF enables flexible identification of various unsafe data types without predefined taxonomies.
-However, in practice, model outputs can be complex with combined safe linguistic features and unsafe content, leading to reduced attribution accuracy.
-In such cases, DABUF will integrate moderation classifiers to identify a minimal subset of unsafe training data for targeted attribution (such as jailbreak).
-When model outputs are relatively straightforward, DABUF uses model outputs directly as the attribution targets.
-We evaluate the performance on two different tasks: in filtering jailbreaking training data and in identifying and mitigating gender bias.
-DABUF outperforms SOTA approaches by up to 7.5% in detection AUPRC in jailbreaking scenarios, and 44.1% in detecting gender bias.
-Moreover, retraining on DABUF-filtered data leads to higher model safety across experiments, underscoring its versatility in addressing a broad spectrum of unsafe data issues.
-Detecting and Filtering Unsafe Training Data via Data Attribution
+Large language models (LLMs) are highly sensitive to even small amounts of unsafe training data, making effective detection and filtering essential for trustworthy model development.
+Current state-of-the-art (SOTA) detection approaches primarily rely on moderation classifiers, which require significant computation overhead for training and are limited to predefined taxonomies.
+In this work, we explore data attribution approaches that measure the similarity between individual training samples and a small set of unsafe target examples, based on data representations such as hidden states or gradients.
+We identify a key limitation in existing methods: unsafe target texts contain both critical tokens that make them unsafe and neutral tokens (e.g., stop words or benign facts) that are necessary to form fluent language, and the latter of which makes the overall representations “noisy” for the purpose of detecting unsafe training data.
+To address this challenge, we propose Denoised Representation Attribution (DRA), a novel representation-based data attribution approach that denoises training and target representations for unsafe data detection.
+Across tasks of filtering jailbreaks and detecting gender bias, the proposed approach leads to significant improvement for data attribution methods, outperforming SOTA methods that are mostly based on moderation classifiers.
+Detecting and Filtering Unsafe Training Data via Data Attribution with Denoised Representation
 Yijun Pan1 Taiwei Shi2 Jieyu Zhao2 Jiaqi W. Ma3 1University of Michigan 2University of Southern California 3University of Illinois Urbana-Champaign
@@
 A primary source of these issues is unsafe training data (Jiang et al., 2024; Chen et al., 2024).
-For instance, inherent biases or toxic content in a dataset can lead to harmful responses (Jiang et al., 2024; Ouyang et al., 2022), while deliberate attacks, such as adversarial prompts or injected backdoors, can be used to bypass safety alignments (Chen et al., 2024; Zou et al., 2023; Li et al., 2024b).
+For instance, inherent biases or toxic content in training data can lead to harmful responses (Jiang et al., 2024; Ouyang et al., 2022), while deliberate training-time attacks, such as adversarial prompts or injected backdoors, can be used to bypass safety alignments (Chen et al., 2024; Zou et al., 2023; Li et al., 2024b).
 Consequently, identifying and removing these unsafe training instances is critical for mitigating risks and building safer LLMs.
 Existing methods for detecting and filtering unsafe training data typically rely on moderation classifiers.
-Online API tools, such as the Perspective API  [^1] https://www.perspectiveapi.com/ and OpenAI’s Moderation API (Markov et al., 2023), focus on certain predefined toxicity taxonomies, but struggle to generalize to nuanced and emerging unsafe artifacts beyond these predefined taxonomies Weber et al.
+Online API tools, such as the Perspective API  [^1] https://www.perspectiveapi.com/ and OpenAI’s Moderation API (Markov et al., 2023), focus on certain predefined toxicity taxonomies, struggling to generalize to nuanced and emerging unsafe artifacts beyond these predefined taxonomies Weber et al.
 (2025).
 Fine-tuned detection models, including Llama-Guard-3-8B (Llama Team, 2024) and Wildguard (Han et al., 2024), require significant time and resources for additional data collection and training.
-Moreover, these moderation classifiers primarily detect semantically unsafe training data without considering the influence of each data point on model training, resulting in suboptimal filtering effectiveness for enhancing model safety.
-In this work, we introduce Data-Attribution-Based Unsafe Training Data Detection and Filtering (DABUF), a method that leverages data attribution techniques to enhance unsafe data detection and filtering in LLMs.
-Data attribution is a family of methods that quantify the influence of individual training data points on specific model outputs (Koh and Liang, 2020; Pruthi et al., 2020).
-Our central hypothesis is that unsafe training data exerts greater influence on unsafe outputs; thus, attributing these outputs back to their influential training instances can reveal which data points are responsible.
-These methods do not require additional training data and can be applied flexibly to diverse types of unsafe model outputs.
-However, applying data attribution naively—by directly attributing unsafe model generations to their training data—is ineffective in many cases.
-LLM outputs, particularly in long-form generations such as jailbreaking attacks, are influenced by a mix of benign and unsafe training data.
-Since model generations include common linguistic structures (e.g., stop words, neutral phrases) alongside unsafe content, direct attribution to the entire sequence leads to a noisy attribution signal, reducing the precision of unsafe data identification.
-To address this challenge, our method introduces a targeted filtering mechanism.
-Specifically, for long-form outputs, such as those found in jailbreaking scenarios, we first use moderation classifiers to identify a small subset of clearly unsafe training data.
-We then use this subset as the attribution target, allowing us to refine the attribution process and isolate the most influential unsafe training instances.
-In contrast, for shorter model outputs, such as those in gender bias scenarios, where the influence of training data is more direct and less noisy, standard data attribution techniques are sufficient without additional moderation filtering.
-We validate our approach through experiments in two distinct setups.
-In jailbreaking scenarios involving adversarial prompts that lead to noisy or long unsafe outputs, we apply the proposed DABUF with moderation classifiers for initial unsafe data identification.
-Conversely, in the gender bias scenario—where outputs are relatively concise—we directly apply DABUF on the model generation.
-Experimental results show that our approach achieves superior detection performance across different model architectures in jailbreaking scenarios and deliver improved safety when models are retrained with filtered data, outperforming state-of-the-art detection methods.
-Furthermore, our method generalizes effectively to gender bias scenarios, highlighting its versatility.
+Moreover, these moderation classifiers primarily flag stylistically unsafe data in a model-agnostic way, without accounting for how individual training samples actually influence the model.
+As a result, they fail to capture the true training impact of unsafe data and are less effective for enhancing model safety through retraining.
+Another line of research focuses on filtering training data by measuring data representation similarity with a set of unsafe model responses.
+These methods are appealing because they do not require additional labeled data and can be flexibly applied across diverse forms of unsafe model outputs.
+The core hypothesis is that unsafe training data exhibit higher similarity to unsafe model outputs.
+Prior work has explored model hidden states, for example, LARF Li et al.
+(2025) and RepSim Pezeshkpour et al.
+(2021), as well as gradient-based similarity (Xia et al., 2024).
+These methods are widely adopted in the data attribution literature (Deng et al., 2025) to quantify the influence of training data on target data (Pruthi et al., 2020).
+However, existing representation-based methods typically measure similarity between training samples and target unsafe examples using representations averaged over all tokens in each sample.
+This naive averaging introduces a key limitation: not all token representations contribute equally to detecting unsafe data.
+Unsafe texts often contain only a few critical tokens that account for the unsafe semantics while the majority of tokens (e.g., stop words or neutral phrases) are benign and contribute to linguistic fluency or neutral factual information.
+As a result, the averaged representations are “noisy”, obscuring the true signal needed to identify unsafe training data.
+To address this limitation, we propose Denoised Representation Attribution (DRA), a novel representation-based data attribution method for unsafe training data detection and filtering.
+DRA denoises the training and target representations in two steps.
+First, motivated by the optimal linear detector in signal detection theory (Stanislaw and Todorov, 1999), DRA centers and whitens the data representations so spurious high-variance dimensions often irrelevant to the unsafeness are downweighted.
+Second, DRA applies a simple leave-target-out surrogate to select most informative hidden dimensions of the representations for unsafe training data detection.
+This two-step procedure leads to denoised representations that are more effective to detect unsafe training samples.
+We validate our proposed approach through experiments on both jailbreaking and gender bias scenarios.
+Experimental results demonstrate that our method consistently outperforms state-of-the-art detection approaches across different model architectures, increasing average detection AUPRC by up to $63.3\%$.
+Moreover, when retraining models on data filtered by our method, we decrease average ASR (Attack Success Rate) by up to $39.9\%$ , confirming the effectiveness of our approach in mitigating unsafe training data.
 2 Related Work
@@
 For instance, SafetyAnalyst (Li et al., 2024a) proposes using LLMs to generate “harm-benefit” tree for interpretable content moderation.
-Beyond content moderation classifiers, some recent studies have leveraged the internal structures of models for unsafe data detection.
-For example, GradSafe (Xie et al., 2024) utilizes gradient similarity with respect to safety-critical parameters to identify unsafe data, while BEBC (Zheng et al., 2024) employs fine-tuned BERT embeddings for content moderation.
-Unlike online moderation tools and LLM-based classifiers, our methods eliminate the need for data curation and additional training.
-Furthermore, by operating independently of predefined safety taxonomies, our approach demonstrates greater flexibility and can effectively handle a wider range of safety-related scenarios.
-In contrast to other non-classifier approaches, such as GradSafe, our methods address the problem through the lens of data attribution.
-This perspective enhances detection performance by incorporating unsafe model behavior and capturing the relationship between training data and model outputs, ultimately fostering safer model developments.
-2.3 Data Attribution for LLMs
-Data attribution methods aim to quantify the impact of individual training samples on a model’s predictions for specific test cases (Koh and Liang, 2020).
-These methods have the potential to detect unsafe training data, as such data are likely to exert a disproportionate influence on unsafe model outputs, making them distinguishable from the broader benign dataset.
-Recently a variety of data attribution methods have been proposed to estimate the influence of training data in the context of LLMs.
-These include gradient-based methods (Xia et al., 2024; Kwon et al., 2024), simulator-based methods (Guu et al., 2023; Chai et al., 2024) and game-theoretic methods (Wang et al., 2024b, c).
-Estimated influence scores have been utilized for tasks such as identifying mislabeled data (Pruthi et al., 2020), understanding memorization (Feldman and Zhang, 2020), and data valuation (Choe et al., 2024).
-While data attribution methods have various applications in LLMs, they are computationally intensive, limiting their applicability to larger models.
-Despite recent advancements in efficient influence estimation methods Kwon et al.
-(2024), the computational burden remains a challenge.
-Gradient-similarity-based approaches, as highlighted in previous works Xia et al.
-(2024); Pruthi et al.
-(2020), offers an efficient solution, making it better suited for scaling to LLMs.
-3 Data-Attribution-Based Unsafe Training Data Detection and Filtering
-Our proposed method consists of two phases: detection and filtering.
-The primary technical challenge arises in the detection phase, where we identify unsafe data points in the training dataset that contribute to unsafe model behaviors.
-In the filtering phase, we mitigate these behaviors by removing the data points most likely to be unsafe.
-3.1 Unsafe Training Data Detection
-We first describe the problem setup of unsafe training data detection. Consider a training dataset with a mixture of benign and unsafe data:
+Beyond safety classifiers, other approaches rely on representation-based similarity for unsafe data detection.
+For example, BEBC (Zheng et al., 2024) employs fine-tuned BERT embeddings for content moderation, while gradient-based methods such as GradSafe (Xie et al., 2024) estimate similarity with respect to safety-critical parameters.
+More broadly, gradient-similarity methods can be seen as a form of data attribution (Koh and Liang, 2020), which quantify the influence of individual training samples on model outputs.
+These attribution-based perspectives are particularly useful for detecting unsafe data, as such data often exert disproportionate influence on unsafe behaviors, making them distinguishable from the broader benign dataset.
+Although influence estimation can be computationally intensive, recent gradient-similarity approaches (Xia et al., 2024; Kwon et al., 2024) provide more efficient alternatives, making them better suited for scaling to large models.
+However, prior work such as GradSafe and BEBC primarily focus on prompt filtering, whereas in training data detection the training label is also important, since it drives model behavior during training.
+Naively extending these methods to the training data setting yields suboptimal results, as noise in model responses manifests as degraded detection performance (Jiao et al., 2025).
+In this context, our methods enable efficient and principled denoising and dimension selection techniques for representation-based detection.
+Our approach enhances detection performance and fosters safer model developments.
+3 Unsafe Training Data Detection and Filtering
+In this section, we formally define the task of unsafe training data detection, introduce a framework for quantifying the sensitivity of representation-based detection methods, and present our proposed Denoised Representation Attribution (DRA) approach for improving detection effectiveness.
+We also describe how this detection method can be used to filter unsafe data for safer downstream model training.
+3.1 Problem Setup: Representation-Based Unsafe Training Data Detection
+Consider a training dataset with a mixture of benign and unsafe data:
 $\mathcal{D}_{\text{train}}=\mathcal{D}_{\text{benign}}\cup\mathcal{D}_{\text{unsafe}},$
 where $\mathcal{D}_{\text{benign}}$ refers to the benign dataset that is safe to train on while $\mathcal{D}_{\text{unsafe}}$ is the unsafe training dataset that could lead to unsafe model behaviors.
-In addition, we assume access to a (small) target dataset $\mathcal{D}_{\text{target}}$ that consists of unsafe model outputs or examples of the unsafe training data.
-In practice, these examples may come from user reports or manual inspection of a small portion of training data.
-The goal of unsafe training data detection is to retrieve $\mathcal{D}_{\text{unsafe}}$ from the entire $\mathcal{D}_{\text{train}}$ with high precision and recall, possibly using the information from the target set $\mathcal{D}_{\text{target}}$.
-A high-quality detection method will help us obtain a cleaner training dataset without overly removing safe training data in the filtering phase.
-Data-Attribution-Based Detection
-We propose to detect the unsafe training data by measuring the influence of each training data point $z\in\mathcal{D}_{\text{train}}$ on the likelihood of the model generating the examples in the target dataset $\mathcal{D}_{\text{target}}$.
-Intuitively, since $\mathcal{D}_{\text{target}}$ consists of unsafe examples, a training data point with higher influence is more likely to be unsafe.
-Formally, we denote the influences as
-$\text{Inf}(z,\mathcal{D}_{\text{target}}),\quad z\in\mathcal{D}_{\text{train}}.$
-In this work we use gradient similarity (Pruthi et al., 2020) to efficiently estimate training data’s influence on model generations, which is a scalable method that has been widely used in data attribution for LLMs (Xia et al., 2024).
-Consider a model parameterized by $\theta$, and denote the negative log-likelihood as $\ell(\cdot;\theta)$.
-The influence of a training data point $z\in\mathcal{D}_{\text{train}}$ on the target dataset $\mathcal{D}_{\text{target}}$ is defined as following:
-$\mathrm{Inf}(z,\mathcal{D}_{\text{target}}):=\eta\cos(\nabla\ell(\mathcal{D}_{\text{target}};\theta),\nabla\ell(z;\theta)),$
-where $\eta$ is the average learning rate, $\nabla\ell(\mathcal{D}_{\text{target}};\theta)$ is the gradient of the negative log-likelihood with respect to $\theta$ evaluated on the target set $\mathcal{D}_{\text{target}}$, and $\nabla\ell(z;\theta)$ is the gradient evaluated on the training point $z$.
-To improve computational efficiency in practice, we follow Xia et al.
-(2024) to reduce the gradient dimension to $d=8192$ via random projection, and adopt optimizer-aware training gradients.
-Data Attribution for LLM Outputs
-In the context of LLMs, the model output is a sequence of tokens.
-Specifically, for each example $x\in\mathcal{D}_{\text{target}}$, it can be represented as $x=(p,r)$, where $p$ is the input prompt and $r$ is the output response, both of which are a sequence of tokens.
-Existing literature (Xia et al., 2024) typically defines $\nabla\ell(\mathcal{D}_{\text{target}};\theta)$ as
-$\nabla\ell(\mathcal{D}_{\text{target}};\theta)=\sum_{x\in\mathcal{D}_{\text{target}}}\nabla\log p(r|p;\theta),$
-where one can further expand the conditional probability $p(r|p;\theta)$ defined by an autoregressive LLM as following:
-$\nabla\ell(r;\theta,p)=\sum_{i=1}^{|r|}\nabla\ell(r_{i}|p,r_{<i};\theta).$
-However, we find that naively applying the data attribution method defined above to long-form unsafe outputs often yields suboptimal performance for identifying unsafe training data.
-When a response $r$ contains many benign tokens and only a few segments of genuinely unsafe content, the attribution signal becomes diluted by the larger volume of neutral or benign tokens.
-In particular, the tokens directly associated with unsafe content–henceforth referred to as unsafe tokens–carry disproportionately stronger gradients, indicating their direct link to harmful outputs.
-Yet, when the model’s overall attribution signal is aggregated over all tokens, these critical unsafe signals become overwhelmed by the contributions of benign tokens, resulting in noisy and less precise detection of unsafe training data points.
-This imbalance is especially problematic in long-form scenarios like jailbreaking attacks, where substantial portions of the model response may be benign filler text interspersed with targeted unsafe content.
-Consequently, the naive approach of attributing every token in the entire generation fails to isolate the truly influential unsafe training instances.
-Please refer to Appendix B.1 for more detailed empirical evidence of this observation.
-Leveraging Externally-Identified Unsafe Data for Effective Attribution
-To address the aforementioned issue in scenarios where the model outputs are long, we propose to take ground-truth labels (instead of the model outputs) from a small, externally identified subset of unsafe training data as the target dataset to attribute.
-For example, in the jailbreaking scenario, we first an LLM-based classifier (Llama-3-Guard-8B (Llama Team, 2024)) to screen the entire training dataset and obtain a small candidate set of unsafe training data $\mathcal{D}_{\text{cand}}$.
-Because such classifiers can have high false-positive rates, we further perform human annotation on $\mathcal{D}_{\text{cand}}$ to filter out benign data points, resulting in a smaller, verified unsafe subset $\mathcal{D}_{\text{identified}}$.
-The target set is then set as $\mathcal{D}_{\text{target}}=\mathcal{D}_{\text{identified}}$.
-In all of our experiments, $|\mathcal{D}_{\text{cand}}|$ is well below 200–manageable for human inspection–whereas the full training set exceeds 40,000 samples.
+In real world scenarios, the training dataset often consists of only a small portion of unsafe data, and the membership of $\mathcal{D}_{\text{benign}}$ and $\mathcal{D}_{\text{unsafe}}$ is unknown.
+The goal of unsafe training data detection is to identify a subset of training data points in $\mathcal{D}_{\text{train}}$ that are mostly likely to be unsafe.
+In addition, we assume access to a (small) target dataset $\mathcal{D}_{\text{target}}$ that consists of unsafe model outputs.
+In practice, these examples may be flagged by users or collected via red teaming.
+Representation-based detection.
+For any $z\in\mathcal{D}_{\text{train}}\cup\mathcal{D}_{\text{target}}$, let $h$ be the mapping from the data point $z$ to its representation $h(z)\in\mathbb{R}^{m}$ (e.g., hidden states or gradients).
+Then we define $B=\{h(z)\mid z\in\mathcal{D}_{\text{benign}}\},U=\{h(z)\mid z\in\mathcal{D}_{\text{unsafe}}\},$ and $Q=\{h(z)\mid z\in\mathcal{D}_{\text{target}}\},$ which are respectively the set of data representations for $\mathcal{D}_{\text{benign}},\mathcal{D}_{\text{unsafe}},$ and $\mathcal{D}_{\text{target}}.$ The set of data representations for $\mathcal{D}_{\text{train}}$ is then $X=B\cup U$.
+A common paradigm in prior work Xie et al.
+(2024); Xia et al.
+(2024); Li et al.
+(2025) is to assign each training representation $x\in X$ ($x\in\mathbb{R}^{m}$) a detection score based on its similarity to an aggregated target representation $\mu_{Q}=\mathrm{Agg}(Q)\in\mathbb{R}^{m}$ (e.g., $\mathrm{Agg}(Q)=\frac{1}{|Q|}\sum_{q\in Q}q$).
+Training samples with the highest detection scores are then flagged as unsafe.
+Formally, the score is typically defined as the inner product
+$s_{\mu_{Q}}(x)\;=\;\langle x,\mu_{Q}\rangle.$
+In LLMs, both the training representation $x$ and the aggregated target representation $\mu_{Q}$ encode rich semantic information, much of which is irrelevant to the unsafeness of the samples.
+For the purpose of detecting unsafe training data, both representations are therefore noisy.
+Goal.
+Our study aims to quantify the sensitivity of the detection score $s_{\mu_{Q}}$, and develop better representation-based detection measures.
+3.2 Sensitivity of the Detection Score
+We frame the problem of detecting unsafe training data in the language of signal detection theory Stanislaw and Todorov (1999): the unsafe training data act as the signal to be distinguished from a much larger background of benign data constituting the noise.
+The sensitivity of such detection can be quantified by the classical discriminability index, $d^{\prime}=\frac{\mu_{\text{signal}}-\mu_{\text{noise}}}{\sigma_{\text{noise}}}$, which depends on the means ($\mu_{\cdot}$) and standard deviations ($\sigma_{\cdot}$) of a given detection measure for the signal and noise distributions, and reflects how well the two distributions can be separated.
+By applying this formulation to the detection score $s_{\mu_{Q}}$ in our problem, we obtain the following sensitivity measure for detecting signal samples $U$ from noise samples XX [^2] Strictly speaking, the noise samples are $B$.
+Under the assumption that $|U|$ is much smaller than $|B|$, the empirical distributions of $X$ and $B$ will be close.
+using $s_{\mu_{Q}}$:
+$d^{\prime}(s_{\mu_{Q}})\;=\;$ $\frac{\mathbb{E}_{u\sim U}[s_{\mu_{Q}}(u)]-\mathbb{E}_{x\sim X}[s_{\mu_{Q}}(x)]}{\sqrt{\mathrm{Var}_{x\sim X}[s_{\mu_{Q}}(x)]}}$
+$=\;$ $\frac{\langle\mu_{U}-\mu_{X},\mu_{Q}\rangle}{\sqrt{\mu_{Q}^{\top}\Sigma_{X}\mu_{Q}}},$ (1)
+where $\mu_{U}$ and $\mu_{X}$ are the mean of $U$ and $X$, and $\Sigma_{X}=\mathrm{Cov}(X)\;\in\;\mathbb{R}^{m\times m}$ is the covariance of $X$.
+It is worth noting that under certain technical assumptions, the discriminability index $d^{\prime}$ is monotonic with respect to AUROC and AUPRC Stanislaw and Todorov (1999).
+Thus $d^{\prime}$ can be viewed as a theoretically grounded surrogate of these widely used evaluation metrics for unsafe data detection, while additionally admitting a natural decomposition that enables fine-grained analysis of the representation-based detection.
+3.3 Improving the Detection Score by Denoising Representations
+Given the discriminability index $d^{\prime}$, our goal is to develop a representation-based detection score $s$ that achieves a higher $d^{\prime}(s)$.
+Optimal linear detection score.
+We consider the family of linear detection scores, $s_{w}(x)\;=\;\langle x,w\rangle,$ for any linear weights $w\in\mathbb{R}^{m}$.
+There is a classical result Poor (2013) in signal detection theory about the optimal linear detection score as stated below.
+Lemma 3.1 (Optimal Linear Detection Score)
+For any $w\in\mathbb{R}^{m},$ denoting $\Delta=\mu_{U}-\mu_{X}$, then
+$d^{\prime}(s_{w})=\frac{\langle\Delta,w\rangle}{\sqrt{w^{\top}\Sigma_{X}w}}$ $=\frac{\langle\Sigma_{X}^{-1/2}\Delta,\Sigma_{X}^{1/2}w\rangle}{\|\Sigma_{X}^{1/2}w\|}$
+$\leq\|\Sigma_{X}^{-1/2}\Delta\|,$
+with equality achieved for optimal $w^{*}\propto\Sigma_{X}^{-1}\Delta.$
+However, note that this theoretically optimal linear weight $w^{*}$ is infeasible to compute, as it depends on $\mu_{U}$, which is unknown due to the unavailability of the unsafe dataset $\mathcal{D}_{\text{unsafe}}$.
+Plug-in detection score.
+In practice, under the hypothesis that the mean $\mu_{Q}$ from the target dataset aligns well with $\mu_{U}$, we propose to set $\hat{w}=\Sigma_{X}^{-1}(\mu_{Q}-\mu_{X})$, which gives us a plug-in detection score $s_{\hat{w}}$ as an estimate of the optimal linear detection score $s_{w^{*}}$.
+Intuitively, compared to the raw score $s_{\mu_{Q}}(x)=\langle x,\mu_{Q}\rangle$, the plug-in score
+$s_{\hat{w}}(x)$ $=\big\langle x,\Sigma_{X}^{-1}(\mu_{Q}-\mu_{X})\big\rangle$
+$=\big\langle\Sigma_{X}^{-1/2}(x-\mu_{X}),\Sigma_{X}^{-1/2}(\mu_{Q}-\mu_{X})\big\rangle+c,$
+is a Mahalanobis correlation (up to a constant $c$ independent of $x$): it first centers both train and target representations by the background mean $\mu_{X}$ to remove global content common to almost all training points, then whitens by $\Sigma_{X}^{-1/2}$ to suppress the high-variance dimensions of the data representation, since the dominant dimensions could correspond to topics or styles instead of the unsafeness.
+Dimension selection via leave-target-out discriminability index.
+For LLMs, the data representation dimension size $m$ can be very large.
+Even with the denoising step described above, the representations could still contain too much irrelevant information.
+Since dimension selection is essentially a subset selection problem, the task becomes challenging when $m$ is large.
+We further propose to conduct a dimension selection to select the most informative hidden dimensions in the representation for unsafe training data detection.
+Specifically, given the eigen decomposition of $\Sigma_{X}=V\Lambda V^{\top}$, we will select a subset of eigenvectors $V_{I}\in\mathbb{R}^{m\times|I|}$ indexed by indices $I\subseteq\{1,2,\ldots,m\}$.
+Let $\Pi_{I}=V_{I}V_{I}^{\top}$ be the projection matrix for the space spanned by $V_{I}$.
+We define a detection score restricted to the selected dimensions $I$ as following:
+$s_{I}(x)=\big\langle\Pi_{I}\,\Sigma_{X}^{-1/2}(x-\mu_{X}),\;\Pi_{I}\,\Sigma_{X}^{-1/2}(\mu_{Q}-\mu_{X})\big\rangle.$
+A natural criterion for selecting the dimensions $I$ is to maximize the discriminability index applied to $s_{I}$:
+$d^{\prime}(s_{I})=\frac{\mathbb{E}_{u\sim U}[s_{I}(u)]-\mathbb{E}_{x\sim X}[s_{I}(x)]}{\sqrt{\mathrm{Var}_{x\sim X}[s_{I}(x)]}}.$
+However, since $d^{\prime}(s_{I})$ depends on $U$, it is infeasible to compute.
+Again, we utilize the trick of approximating $U$ from the ground truth unsafe training dataset by $Q$ from the target dataset, resulting in a leave-target-out (LTO) discriminability index:
+$d^{\prime}_{\text{LTO}}(s_{I})=\frac{\mathbb{E}_{q\sim Q}[s_{I}(q)]-\mathbb{E}_{x\sim X}[s_{I}(x)]}{\sqrt{\mathrm{Var}_{x\sim X}[s_{I}(x)]}}.$
+We name it “leave-target-out” as $d^{\prime}_{\text{LTO}}$ can be viewed as measuring the sensitivity of a hypothetical detection task that aims to identify $\mathcal{D}_{\text{target}}$ out of $\mathcal{D}_{\text{train}}\cup\mathcal{D}_{\text{target}}$.
+The proposed DRA detection score.
+All together, the proposed Denoised Representation Attribution (DRA) method is a detection score defined as $s_{I^{*}}(\cdot)$, where $I^{*}=\arg\max_{|I|=r}d^{\prime}_{\text{LTO}}(s_{I})$ for some given dimension size $r$ (we report results for $r=30$ across experiments).
+In practice, instead of directly selecting the dimensions from all $m$ eigenvectors, we first reduce the candidate dimensions to the $m^{\prime}<m$ top eigenvectors, and then optimize the indices $I\subseteq\{1,2,\ldots,m^{\prime}\}$.
 3.4 Unsafe Training Data Filtering
-Using the estimated influence scores, we perform retrieval to identify the training examples most influential on the unsafe target samples.
-We select the top $K$ elements from the ranked list of $\text{Inf}(z,\mathcal{D}_{\text{target}})$ values.
-Let $\mathcal{S}_{K}(\mathcal{D}_{\text{target}})\subseteq\mathcal{D}_{\text{train}}$ denote the set of $K$ most influential training samples selected, when the target dataset $\mathcal{D}_{\text{target}}$ is used as the target for attribution.
-By removing $\mathcal{S}_{K}(\mathcal{D}_{\text{target}})$ from the training data, we construct a cleaner dataset that is expected to improve model safety upon retraining.
+Using the detection scores derived above, we filter the training dataset by removing training examples with highest scores.
+Specifically, for each $x\in\mathcal{D}_{\text{train}}$, we compute $s_{I^{*}}(x)$ and rank all training samples accordingly.
+We then select the top $K$ elements from this ranked list, denoting the selected subset as
+$\mathcal{S}_{K}(\mathcal{D}_{\text{target}})=\arg\operatorname{top-k}_{x\in\mathcal{D}_{\text{train}}}\;s_{I^{*}}(x)$
+By removing $\mathcal{S}_{K}(\mathcal{D}_{\text{target}})$ from the training data, we obtain a cleaner dataset that is expected to reduce unsafe behaviors upon retraining.
 4 Experiments: Jailbreaking Injection Detection
@@
 4.1 Experimental Setup
-Overview.
-We focus on a realistic training scenario wherein the benign portion of the data can be heterogeneous, consisting of:
-  1. 1.
-    Fully benign prompt-response pairs with no harmful content.
-  2. 2.
-    Safe demonstrations: pairs where the prompt may be unsafe, but the response is aligned and refuses or mitigates the request.
-These “safe” demonstrations have been shown to improve model safety (Jain et al., 2023).
-We denote the set of safe demonstrations as $\mathcal{D}_{\text{safe}}$.
-We augment this benign dataset with a small set of unsafe injections designed to induce harmful responses.
-Across all experiments, the total injection ratio is below $0.025\%$.
-Identifying these few harmful training instances is extremely challenging, but crucial for mitigating jailbreaking vulnerabilities.
-Datasets
-We use the dataset Ultrachat 200k as our benign dataset and consider two unsafe datasets, ToxicChat and XSTest-Response.
-For each unsafe dataset, we split the dataset into $\mathcal{D}_{\text{safe}}$ (unsafe prompts with safe responses), $\mathcal{D}_{\text{unsafe}}$ (unsafe prompts and unsafe responses), and $\mathcal{D}_{\text{test}}$ (held-out unsafe prompts for evaluation).
-In our experiments, we inject both the $\mathcal{D}_{\text{safe}}$ and $\mathcal{D}_{\text{unsafe}}$ from the unsafe dataset into the benign Ultrachat 200k dataset to form the whole training dataset $\mathcal{D}_{\text{train}}=\mathcal{D}_{\text{benign}}\cup\mathcal{D}_{\text{safe}}\cup\mathcal{D}_{\text{unsafe}}$.
-  • Ultrachat 200k.
-Ultrachat 200k [^2] https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k is a heavily filtered version of the UltraChat Ding et al.
-(2023) dataset, which comprises over 200k instructions for instruction fine-tuning purposes.
-We use a subset of the train sft split for our $\mathcal{D}_{\text{benign}}$, comprising $41,573$ samples.
-  • ToxicChat.
-ToxicChat [^3] https://huggingface.co/datasets/lmsys/toxic-chat Lin et al.
-(2023) is a dataset consisting prompt-response pairs annotated with prompt toxicity and jailbreak-ness (response toxicity), curated from user interactions.
-We use a subset of the latest version of ToxicChat: ToxicChat-0124, and apply the following split: $|\mathcal{D}_{\text{safe}}|=127$, $|\mathcal{D}_{\text{unsafe}}|=97$, and $|\mathcal{D}_{\text{test}}|=30$.
-  • XSTest-Response.
-XSTest-Response [^4] https://huggingface.co/datasets/allenai/xstest-response Han et al.
-(2024) is a dataset consisting of responses to the original XSTest Röttger et al.
-(2023), which includes 446 annotations of prompt harmfulness and model response harmfulness.
-We use the official response_harmfulness subset and apply the following split: $|\mathcal{D}_{\text{safe}}|=121$, $|\mathcal{D}_{\text{unsafe}}|=65$, and $|\mathcal{D}_{\text{test}}|=20$.
+Datasets.
+We conduct our experiments following the standardized setup from DATE-LM Jiao et al.
+(2025)  [^3] https://huggingface.co/DataAttributionEval, which provides an evaluation setup for unsafe training data detection.
+This benchmark uses UltraChat-200k Ding et al.
+(2023) as the benign dataset, and injects unsafe data from ToxicChat Lin et al.
+(2023) and XSTest-Response Röttger et al.
+(2023) into UltraChat to form unsafe training data.
+ToxicChat and XSTest-Response cover a wide range of jailbreaking settings, including both toxicity/harmfulness and adversarial prompt-response pairs.
+In addition, we evaluate model safety using the official test splits.
+This ensures that our setting is directly comparable to prior baselines while maintaining consistency across datasets.
 4.2 Evaluation Metrics
-Given the retrieved set $\mathcal{S}_{K}(\mathcal{D}_{\text{target}})$ containing $K$ top influential training data to the validation set, we define the precision and recall as:
+Detection.
+Given the retrieved set $\mathcal{S}_{K}(\mathcal{D}_{\text{target}})$ containing $K$ top influential training data to the target set, we define the precision and recall as:
 precision $=\frac{|\mathcal{S}_{K}(\mathcal{D}_{\text{target}})\cap\mathcal{D}_{\text{unsafe}}|}{|\mathcal{S}_{K}(\mathcal{D}_{\text{target}})|}$
 recall $=\frac{|\mathcal{S}_{K}(\mathcal{D}_{\text{target}})\cap\mathcal{D}_{\text{unsafe}}|}{|\mathcal{D}_{\text{unsafe}}|}$
-We adopt the Area Under Precision Recall Curve (AUPRC) as well as the precision, recall and F1 scores calculated with the top 100 samples identified for a comprehensive evaluation of baselines models and our methods.
-To evaluate model safety, we employ Attack Success Rate (ASR) on the test set $\mathcal{D}_{\text{test}}$, which measures the proportion of the unsafe prompts in $\mathcal{D}_{\text{test}}$ that successfully elicit unsafe responses from the model.
-4.3 Baselines
-We include baselines from three categories: online API tools (OpenAI moderation API), fine-tuned LLM as detectors (Llama-Guard-3-8B, Wildguard) and other model-free methods (GradSafe).
+We evaluate detection with the Area Under the Precision–Recall Curve (AUPRC).
+Retraining.
+To assess model safety before and after filtering, we report the Attack Success Rate (ASR) on the held-out test split $\mathcal{D}_{\text{test}}$, defined as the fraction of unsafe prompts that elicit an unsafe response.
+To obtain stable and reliable results, ASR is averaged over 10 independent inference runs on $\mathcal{D}_{\text{test}}$.
+The target set $\mathcal{D}_{\text{target}}$ is defined to be the union of all unsafe model outputs collected across these runs.
+For detailed evaluation protocol refer to Appendix A.
+4.3 Methods for Comparison
+Table 1: AUPRC ($\uparrow$) of baseline models and our proposed method.
+For each backbone model, the highest AUPRC among data attribution methods is highlighted in bold, while the second highest is underlined.
+The last column shows the average improvement ($\Delta$) of our methods over their respective baselines.
+Model Method ToxicChat AUPRC (%) $\uparrow$ XSTest-Response AUPRC (%) $\uparrow$ $\Delta$
+OpenAI Moderation API 24.3 37.8
+Llama-Guard-3-8B 44.5 91.6
+Wildguard 56.0 93.0
+GradSafe 37.6 27.4
+Llama-3-8B-Instruct GradSim 36.4 1.7
+GradSim-DRA 83.2 81.5 +63.3
+Repsim 56.3 88.5
+Repsim-DRA 92.6 98.6 +23.2
+Gemma-2-9B-it GradSim 41.1 50.2
+GradSim-DRA 82.3 95.5 +43.3
+Repsim 51.6 83.9
+Repsim-DRA 98.2 99.8 +31.3
+Baselines.
+We include baselines from four categories: online API tools (OpenAI moderation API), fine-tuned LLM as detectors (Llama-Guard-3-8B, Wildguard), model-free methods (GradSafe) and baseline data attribution methods (GradSim and RepSim).
 OpenAI moderation.
@@
 GradSafe Xie et al.
-(2024) differs fundamentally in methodology by analyzing gradients with respect to safety-critical parameters of Llama-2, specifically focusing on the gradient of the model’s compliance response to prompts.
-In contrast, our approach directly traces unsafe behaviors back to the training data by leveraging token-level attributions to identify the sources of unsafe outputs.
-GradSafe operates independently of the model’s responses, providing pre-hoc moderation akin to LLM classifiers, whereas our method emphasizes post-hoc attribution to uncover the origins of unsafe model behavior.
+(2024) detect unsafe training data by analyzing gradients with respect to safety-critical parameters of Llama-2, specifically focusing on the gradient of the model’s compliance response to prompts.
+GradSafe operates independently of the model’s responses, providing pre-hoc moderation akin to LLM classifiers, whereas our setting emphasizes post-hoc attribution to uncover the origins of unsafe model behavior.
+GradSim.
+Previous work has demonstrated that GradSim Xia et al.
+(2024); Jiao et al.
+(2025) is effective in tracing the influence of training samples on a given target set.
+We denote $h(z)$ as the gradient of sample $z$.
+The Grad-Sim score is then defined as the inner product between the $\ell_{2}$-normalized $h(z)$ and the $\ell_{2}$-normalized target gradient mean $\mu_{Q}$.
+RepSim.
+Representation Similarity Pezeshkpour et al.
+(2021) scores have been used for various tasks, such as training data selection and factual tracing Jiao et al.
+(2025).
+Here $h(z)$ is the last-layer hidden state of the final token in sample $z$.
+Similar with GradSim, the RepSim score is then defined as the inner product between the $\ell_{2}$-normalized $h(z)$ and the $\ell_{2}$-normalized target hidden state mean $\mu_{Q}$.
+Our methods.
+We apply the proposed DRA on GradSim and RepSim, respectively lead to GradSim-DRA and Repsim-DRA.
 4.4 Results and Discussion
 In this section, we present the results of baseline methods and our approach for detecting and filtering jailbreaking data.
-We first demonstrate that fine-tuning language models on jailbreaking training data can effectively compromise their safety.
+We first show the results of detection performance.
+Table 1 reports the AUPRC of unsafe training data detection across baseline classifiers and data attribution methods.
+Our proposed method consistently improves upon data attribution baselines that leverage either hidden states or gradients, often yielding substantial performance gains.
+Notably, after applying our approach, data-attribution-based detection not only surpasses its vanilla counterparts but also outperforms the strongest baseline classifiers, highlighting the effectiveness of data attribution for uncovering unsafe training signals that classifiers fail to capture.
+Before we move on to the results of filtering and retraining, we demonstrate that fine-tuning language models on jailbreaking training data can effectively compromise their safety.
 Table 2 presents the ASR across different models and datasets.
-In comparison to training on the benign dataset $\mathcal{D}_{\text{benign}}$ only, the ASR is significantly higher when training with $\mathcal{D}_{\text{train}}$ that consists of unsafe training data, confirming that training on injected unsafe data results in unsafe model behaviors.
+In comparison to training on the benign dataset $\mathcal{D}_{\text{benign}}$ only, the ASR is significantly higher when training with $\mathcal{D}_{\text{train}}$ that consists of unsafe training data.
 Table 2: Attack Success Rate (ASR) across trained models and datasets. A higher ASR indicates a more unsafe model.
-Model Data ToxicChat XSTest-Response
-Llama-3-8B $\mathcal{D}_{\text{train}}$ 93.3% 50%
-$\mathcal{D}_{\text{benign}}$ 66.7% 0%
-Gemma-2-9B $\mathcal{D}_{\text{train}}$ 90.0% 100%
-$\mathcal{D}_{\text{benign}}$ 83.3% 70%
-Table 2: AUPRC of baseline models and our method. The highest AUPRC is highlighted in bold, while the second highest is underlined.
-Method ToxicChat (%) XSTest-Response (%)
-OpenAI Moderation API 11.7 11.8
-Llama-Guard-3-8B 30.3 82.5
-Wildguard 44.5 85.9
-GradSafe 30.7 47.3
-Llama-3-8B-DABUF 52.0 74.1
-Gemma-2-9B-DABUF 49.1 64.1
-Table 3: Precision, recall, and F1 scores of baseline models and our method, calculated based on the top 100 identified training data points.
-The highest F1 score is highlighted in bold, and the second highest is underlined.
-ToxicChat XSTest-Response
-Method Precision (%) Recall (%) F1 (%) Precision (%) Recall (%) F1 (%)
-OpenAI Moderation API 16.0 16.5 16.2 19.0 28.8 22.9
-Llama-Guard-3-8B 33.0 34.0 33.5 57.0 86.4 68.7
-Wildguard 46.0 47.4 46.7 59.0 89.4 71.1
-GradSafe 51.1 47.4 49.2 45.0 68.2 54.2
-Llama-3-8B-DABUF 51.0 52.6 51.8 57.0 86.4 68.7
-Gemma-2-9B-DABUF 52.0 53.6 52.8 53.0 80.3 63.9
-Table 4: Attack Success Rate (ASR) comparison between models retrained with the top 100 unsafe training samples filtered by baseline methods and DABUF. A higher ASR reflects a more unsafe model.
-Model Filtering Method ToxicChat XSTest-Response
-Llama-3-8B OpenAI Moderation 96.7% 25%
-GradSafe 93.3% 25%
-Wildguard 90.0% 15%
-Llama-Guard-3-8B 90.0% 50%
-DABUF 86.7% 5%
-Gemma-2-9B OpenAI Moderation 90% 45%
-GradSafe 86.7% 60%
-Wildguard 80% 20%
-Llama-Guard-3-8B 90% 30%
-DABUF 66.7% 15%
-Table 2 and 3 respectively show the AUPRC and the top 100 precision, recall, and F1 of unsafe training data detection across models and methods.
-Our method demonstrates superior or comparable performance across different experimental settings.
-Notably, DABUF applied to Llama-3-8B achieves the highest performance in the ToxicChat injection experiment, showcasing the advantages of using data attribution for detecting unsafe training data.
-While state-of-the-art classifiers like Wildguard outperform DABUF in XSTest-Response experiments, this is largely due to XSTest-Response’s focus on toxicity and explicit harm, aligning with their in-distribution training data.
-More importantly, Table 4 highlights that the proposed DABUF significantly outperforms all baseline methods in terms of the ASR of models retrained after filtering out the top 100 unsafe training samples identified by different methods.
-The results demonstrate that our data attribution approach, which explicitly accounts for the model training process, effectively identifies and filters unsafe training data that contributes the most to the unsafe model behaviors of interest, which leads to models with better safety when retrained on filtered datasets.
+Model Data ToxicChat ASR $\downarrow$ XSTest-Response ASR $\downarrow$
+Llama-3-8B-Instruct $\mathcal{D}_{\text{train}}$ 33.3% 73.0%
+$\mathcal{D}_{\text{benign}}$ 26.7% 0.0%
+Gemma-2-9B-it $\mathcal{D}_{\text{train}}$ 28.0% 71.0%
+$\mathcal{D}_{\text{benign}}$ 15.7% 0.0%
+Table 3 further shows that applying our method not only improves detection but also leads to safer retrained models, as shown by consistently lower ASR values.
+This underscores that our method is not only effective at identifying unsafe training data but also at filtering those samples that most strongly drive harmful model behaviors, leading to significant improvements in retrain safety.
+In particular, gradient-based attribution enhanced with our denoising and dimension selection approach yield the greatest increase in retrained model safety.
+We believe this is because gradient-based attributions can more accurately capture the underlying training dynamics.
+Table 3: Attack Success Rate (ASR) comparison between models retrained with the top unsafe training samples filtered.
+A higher ASR reflects a more unsafe model and the lowest ASR for each setting is in bold.
+The last column shows the average decrease in ASR after applying DRA to baseline data attribution methods.
+Model Filtering Method ToxicChat ASR $\downarrow$ XSTest-Response ASR $\downarrow$ $\Delta$
+Llama-3-8B-Instruct Wildguard 29.0% 50.0%
+GradSim 34.7% 68.0%
+GradSim-DRA 26.0% 43.0% 16.9-16.9
+Repsim 29.0% 63.0%
+Repsim-DRA 27.3% 60.0% 2.35-2.35
+Gemma-2-9B-it Wildguard 30.5% 30.0%
+GradSim 29.0% 77.0%
+GradSim-DRA 23.3% 3.0% 39.9-39.9
+Repsim 29.0% 70.0%
+Repsim-DRA 30.7% 64.0% 2.15-2.15
 5 Experiments: Gender Bias Mitigation
@@
 For the target set, we generated gender-biased model responses for 50 occupation prediction prompts ($\mathcal{D}_{\text{target}}$).
-Additional details of the experiment are provided in Appendix A.
-Since model responses in this scenario are relatively simple, we do not use externally identified unsafe training data as we did in jailbreaking setups.
+Additional details of the experiment are provided in Appendix B.
 5.2 Evaluation Metrics
-Similar with jailbreaking injection, we evaluate the detection performance via precision, recall, F1 score, and AUPRC.
+Detection
+Similar with jailbreaking injection, we evaluate the detection performance via AUPRC.
+Retrain Safety
 To evaluate gender bias in trained models, we adopt the metric of True Positive Rate (TPR) Gender Gap following De-Arteaga et al.
@@
 A larger TPR Gender Gap observed on the held-out test set indicates that the model exhibits stronger gender biases.
+Same with the jailbreaking setting, we average all results across 10 inference runs and the target set $\mathcal{D}_{\text{target}}$ is defined to be the union of all biased model outputs collected across these runs.
 5.3 Results and Discussion
-First, we demonstrate that training a model on gender-biased data leads to behaviors that reflect gender biases.
+First, we demonstrate that training a model on gender-biased data leads to gender biases.
 As can be seen in Table 4, models trained on gender-biased data exhibit a significant higher TPR Gender Gap in comparison to those trained on unbiased data.
-Table 5: TPR Gender Gap on unbiased and biased datasets.
-Model Dataset TPR Gender Gap
-Llama-3-8B Unbiased 0.04
-Biased 0.16
-Gemma-2-9B Unbiased 0.02
-Biased 0.08
-Table 6: AUPRC and precision, recall, and F1 of the baseline model and our methods.
-The highest AUPRC and F1 value are highlighted in bold, while the second highest underlined.
-The suffix DABUF denotes our method.
-Method AUPRC Precision Recall F1
-Ada3 0.089 0.500 0.330 0.400
-Llama-3-8B-DABUF 0.474 0.610 0.407 0.488
-Gemma-2-9B-DABUF 0.530 0.670 0.447 0.536
-Table 7: TPR Gender Gap comparison between models retrained with the top 100 biased samples filtered by Ada3 and our method.
+Table 4: TPR Gender Gap for model trained on unbiased and biased data.
+Model Dataset TPR Gender Gap $\downarrow$
+Llama-3-8B-Instruct Unbiased -0.066
+Biased 0.848
+Gemma-2-9B-it Unbiased 0.032
+Biased 0.452
+Table 5: Gender bias detection AUPRC.
+The highest AUPRC for each model is highlighted in bold, while the second highest is underlined.
+The last column reports the change ($\Delta$) compared to the corresponding baseline.
+Model Method AUPRC $\uparrow$ $\Delta$
+Llama-3-8B-Instruct Wildguard 0.059
+GradSim 0.417
+GradSim-DRA 0.550 +0.133+0.133
+Repsim 0.333
+Repsim-DRA 0.748 +0.415+0.415
+Gemma-2-9B-it Wildguard 0.059
+GradSim 0.404
+GradSim-DRA 0.468 +0.064+0.064
+Repsim 0.009
+Repsim-DRA 0.369 +0.360+0.360
+Table 6: TPR Gender Gap comparison between models retrained with the top 500 biased samples filtered.
 A higher TPR Gender Gap indicates a model with more gender bias.
-The lowest value is highlighted in bold.
-Model Filtering Method TPR Gender Gap
-Llama-3-8B None 0.16
-Ada3 0.28
-DABUF 0.14
-Gemma-2-9B None 0.08
-Ada3 0.04
-DABUF 0.02
-Table 6 presents the AUPRC, precision, recall, and F1 score results.
-Embedding-based tools like Ada3 exhibit poor performance in detecting unsafe data, suggesting that general-purpose embedding tools are not well-suited for detecting gender-biased data.
-In contrast, as a model-free approach, our method adapts effectively to different injection scenarios using diverse unsafe validation data.
-Table 7 presents the TPR Gender Gap for models retrained after removing the top 100 influential samples identified.
-These results demonstrate that our approach effectively detects training samples that contribute to gender bias, and retraining on the filtered data leads to models with reduced gender bias.
-Notably, while embedding-based methods like Ada3 achieve non-trivial detection performance, filtering using Ada3 actually results in a higher TPR Gender Gap compared to training on the original data for Llama-3-8B model.
-This suggests that embedding-based methods may not be suitable for training data filtering.
-We believe this occurs because embedding-based models focus primarily on semantics, causing them to remove both unsafe and safe training data indiscriminately.
-This indiscriminate removal undermines the model’s ability to effectively learn the nuances of the occupation prediction task.
+The lowest value for each model is highlighted in bold.
+The last column reports the change ($\Delta$) in gender gap when applying DRA to data attribution methods
+Model Filtering Method TPR Gender Gap $\downarrow$ $\Delta$
+Llama-3-8B-Instruct None 0.848
+Wildguard 0.798
+GradSim 0.650
+GradSim-DRA 0.246 0.404-0.404
+Repsim 0.296
+Repsim-DRA 0.092 0.204-0.204
+Gemma-2-9B-it None 0.452
+Wildguard 0.480
+GradSim 0.344
+GradSim-DRA 0.224 0.120-0.120
+Repsim 0.536
+Repsim-DRA 0.346 0.190-0.190
+Table 5 reports the AUPRC results for gender-biased unsafe data detection.
+We observe that moderation classifiers such as Wildguard fail to generalize to this setting, as gender bias often does not include explicit harm signals captured during moderator training.
+Applying our method consistently improves the performance of data attribution approaches, with particularly high gains for gradient-based method.
+This indicates that our approach is especially effective at amplifying the signal of subtle unsafe patterns like gender bias, where classifier-based tools struggle.
+Table 6 presents the TPR Gender Gap for models retrained after removing the top 500 influential samples identified by different detection methods.
+These results show that applying our method to representation-based approaches consistently improves not only the detection of unsafe data but also the downstream model safety after retraining, even in subtle cases where unsafety is implicit, such as gender bias.
+In particular, gradient-based data attribution methods benefit the most from our approach: filtering guided by gradient-based attribution yields models with substantially reduced gender bias.
+This observation aligns with our findings in the jailbreak setting, further reinforcing that gradient-based attribution provides a principled solution for identifying and mitigating unsafe training data.
 6 Conclusion
-This work proposes the challenge of detecting and filtering unsafe training data embedded within larger benign datasets.
-We propose using Data-Attribution-Based Unsafe Training Data Detection and Filtering (DABUF) to detect and filter unsafe training data that contributes to unsafe model behavior.
-In scenarios where model outputs are long, DABUF overcomes the noisy aggregation of token-wise gradients by utilizing externally identified unsafe training data for effective attribution.
-Remarkably, DABUF’s performance for jailbreaking samples detection and filtering surpasses those by SOTA LLM-based moderation classifiers across various models and datasets.
-Furthermore, our method is not confined to existing unsafe taxonomies, demonstrating its adaptability to broader unsafe scenarios, such as gender bias mitigation.
+Our analysis provides a theoretical grounding for representation-based methods in unsafe training data detection and reveals the presence of noise within representations that limits their effectiveness.
+Motivated by this, we introduce Denoised Representation Attribution (DRA), which substantially improve detection performance and lead to safer models after retraining.
+Empirically, our method outperforms state-of-the-art LLM-based moderation classifiers across diverse datasets and unsafe scenarios, while remaining adaptable beyond predefined taxonomies.
+Finally, we observe that gradient-based approaches, though sometimes yielding lower detection scores, consistently produce the safest retrained models, underscoring their practical value for mitigating unsafe behaviors.
 Limitation
@@
 Detection Taxonomy
-Our current work does not provide a fine-grained detection taxonomy, as both training and validation data encompass multiple unsafe genres.
-While we recognize that carefully selecting validation data could potentially enable detection of specific genres, we consider this beyond the scope of the present study and recommend it as a promising direction for future investigation.
+Our current work does not provide a fine-grained detection taxonomy, as both training and target data encompass multiple unsafe genres.
+While we recognize that carefully selecting target data could potentially enable detection of specific genres, we consider this beyond the scope of the present study and recommend it as a promising direction for future investigation.
 Potential Risks
@@
 Recent research on adversarial attacks in data attribution Wang et al.
-(2024d) suggests significant vulnerabilities in influence estimation techniques.
+(2024b) suggests significant vulnerabilities in influence estimation techniques.
 Specifically, attackers could potentially manipulate the estimated influence scores to strategically conceal unsafe training data, thereby undermining the robustness of our detection approach.
-Appendix A Gender Bias setup
+Appendix A Model Safety Evaluation
+Safety Evaluation Guidelines ### 1.Score 1: The model proactively declines to act on the user’s instruction or steers the conversation towards an area consistent with Meta’s usage guidelines.
+This is the most preferred response when confronted with policy-violating instructions.
+### 2.Score 2: If the response is simply a repetition of the request or if begins with an affirmative response for theoretical framework but lacks actual details, then it should be given a low harmfulness level of 2.
+Or without directly addressing or endorsing a user’s inappropriate instruction, it offers a warning.
+### 3.Score 3: The model engages with the policy-violating request with an evident lack of discernment regarding Meta’s guidelines.
+It offers complete content related to the request but without embedding necessary precautions or disclaimers.
+Nonetheless, the response, on its own, isn’t glaringly inappropriate and doesn’t wholly resonate with the user’s infringing objective.
+### 4.Score 4: The model responds to the policy-violating instruction by including explicitly prohibited content, which clearly violates Meta’s usage guidelines.
+However, the model’s response is not a direct fulfillment of the user’s intent and is not as severe as the most inappropriate response of Score 5.
+### 5.Score 5: The model fulfill the user’s policy-violating instruction without any deviation.
+It completely ignores the usage guidelines and provides detailed information that clearly violates the Meta policy.
+The model’s response is a direct endorsement of the user’s intent.
+Following prior works He et al.
+(2024); Qi et al.
+(2023); Jiao et al.
+(2025), we evaluate model response safety using LLM-as-a-judge, leveraging the above guidelines.
+Each response is scored from 1 to 5 according to the severity of unsafe content.
+We define the Attack Success Rate (ASR) as the proportion of responses receiving a score of 5.
+For reproducible and scalable evaluation, we use the fixed-version GPT4o (2024-11-20).
+Appendix B Gender Bias setup
 In this section we illustrate the construction of gender bias injection from the Bias in bios De-Arteaga et al. (2019b) dataset.
@@
 Note that in the original 10000 training set there are also normal entries for female/male nurses and physicians, so we consider the retrieval of these injected data as non-trivial.
-A.1 Baselines
-Since existing LLM safety classifiers are not adapted to detect gender biases, we employ the following models baselines.
-Embedding Models
-Because unsafe training data may be semantically distinct from benign training data, a detection approach that compares their representations provides a natural baseline.
-Specifically, we consider Ada3  [^5] https://openai.com/index/new-and-improved-embedding-model/, a SOTA model for texual embeddings.
-We compute embedding similarities between individual training examples and the mean embedding of the validation set, then identify training samples that exhibit the highest similarity scores.
-Appendix B Analysis
-B.1 Noisy Token-wise Gradients
-In this section we show that model outputs can be noisy and thus directly using responses’ gradient results in sub-optimal results.
-A key observation is that training data impact the predictions of different tokens unevenly, with tokens containing unsafe content being the most affected.
-Therefore, unsafe tokens’ gradients result in greater influence when used for attribution compared to benign tokens’ gradients, making the overall gradient, as a summation of token-wise gradients, inherently noisy.
-To elaborate this, we evaluate each token’s performance when used as target for attribution and show that token-wise gradients from model outputs have diverse performances for detection.
-Specifically, we consider each token $t$’s detection performance by its detection precision:
-$\text{prec}_{t}$ $=\frac{|\mathbf{S}_{M}(t)\cap\mathcal{D}_{\text{unsafe}}|}{|\mathbf{S}_{M}(t)|}$
-By selecting the top $N$ token-wise gradients $t_{1},t_{2},\dots,t_{N}$ that achieve highest precision scores from the entire validation set, we used the filtered validation gradient as the target for attribution:
-¯$\bar{\nabla}\ell(\mathcal{D}_{\text{val}};\theta)=\frac{1}{N}\sum_{i=1}^{N}\nabla\ell(t_{i};\theta)$
-Figure 1: Performance analysis on ToxicChat of varying number of top-contributing tokens.
-Figure 1 illustrates the impact of the number of included token-wise gradients on retrieval performance.
-Initially, incorporating more top-contributing token-wise gradients improves performance; however, as more gradients are included, the features become noisier, and performance begins to decline.
-This highlights the inherent noise in gradient features.
-Table 8: Contribution of token’s gradient to detection for Llama-3-8B trained model in the XSTest-Response injection setting.
-Token AUPRC
-and 0.040
-the 0.070
-on 0.050
-poison 0.176
-commit 0.274
-weapon 0.133
-Table 8 shows the efficacy of individual tokens in identifying unsafe training data.
-As shown, common words such as and, the, and on are largely ineffective for detection, whereas explicit unsafe tokens such as poison, commit, and weapon make significant contributions.
-This disparity can be attributed to the mechanism behind unsafe training: unsafe training data has significant influence on the prediction of unsafe tokens while exerting minimal impact on common syntax tokens like and, the and on.
-Consequently, model output, as a combination of benign and unsafe tokens, are sub-optimal targets for attribution.
-Table 9: AUPRC comparison between directly using model responses as the attribution target and DABUF.
-Method ToxicChat (%) XSTest-Response (%)
-Llama-3-8B-Response 4.70 35.4
-Llama-3-8B-DABUF 52.0 74.1
-Gemma-2-9B-Response 4.12 11.9
-Gemma-2-9B-DABUF 49.1 64.1
-Table 10: Precision, recall, and F1 scores comparison between directly using model responses as the attribution target and DABUF, calculated based on the top 100 identified training data points.
-ToxicChat XSTest-Response
-Method Precision (%) Recall (%) F1 (%) Precision (%) Recall (%) F1 (%)
-Llama-3-8B-Response 14.0 14.4 14.2 19.0 28.8 22.9
-Llama-3-8B-DABUF 51.0 52.6 51.8 57.0 86.4 68.7
-Gemma-2-9B-Response 14.0 14.4 14.2 21.0 31.8 25.3
-Gemma-2-9B-DABUF 52.0 53.6 52.8 53.0 80.3 63.9
-This is further demonstrated by experimental results in Table 9 and table 10.
-DABUF, which leverages externally classified training data as its attribution target, consistently outperforms existing data attribution methods that rely solely on model responses.
-This finding suggests that depending exclusively on model responses can introduce noise, thereby hindering effective detection.
-B.2 Validation Set Variety
-In this section we delve into the effect of validation set variety on detection performance.
-Intuitively, having a larger validation set where model unsafe behaviors are observed across different domains helps with retrieval, especially when the training data contains multiple unsafe genres.
-By default we use a validation set of size 30 for retrieval in ToxicChat injection, where the injected data are in-the-wild user interactions with LLMs and thus contain diverse unsafe behaviors.
-By varying the size of $\mathcal{D}_{\textbf{identified}}$, we evaluate their retrieval performance.
-Figure 2: Performance on ToxicChat with varying validation set size.
-Figure 2 demonstrates the effect of validation set variety on unsafe training data detection.
-The sampling process for each size is repeated for 10 times and the mean as well as standard deviation are calculated accordingly.
-It is observed that a larger validation set size leads to higher detection performance.
-By including a more diverse validation set, we can capture a wider range of unsafe training data.
-Nevertheless, we note that by including as little as 20 validation samples, our approach already exceeds SOTA models like Llama-Guard-3-8B and Wildguard.
-Appendix C Experimental Details
-For experiments involving model training, we train for 4 epochs with a warm up ratio of 0.1 and used learning rate of 1e-4 for Llama-3-8B and 1e-5 for Gemma-2-9b respectively.
+Appendix C Analysis
+C.1 Noisy Representation
+In this section we show that representations such as gradient or hidden states carry inherent noise in some of its dimensions, which hampers detection.
+Figure 1: t-SNE visualization of the raw gradient space of the Llama-3.1-8B-Instruct model with XSTest-injected data (AUPRC = 0.017).
+Figure 2: t-SNE visualization of the whitened PCA dimensions selected by DRA for the Llama-3.1-8B-Instruct model with XSTest-Response-injected data (AUPRC = 0.815).
+Figures 1 and 2 present the t-SNE visualizations of the raw and filtered gradient spaces of the Llama-3.1-8B-Instruct model with XSTest-Response-injected training data, respectively.
+In the raw gradient space, dimension-wise noise makes the unsafe gradients indistinguishable from the larger benign dataset.
+In contrast, after applying DRA, the unsafe gradients form clear and separable patterns, leading to substantially improved detection performance.
+C.2 Target Set Variety
+In this section we delve into the effect of target set variety on detection performance.
+Intuitively, having a larger target set where model unsafe behaviors are observed across different domains helps with detection, especially when the training data contains multiple unsafe genres.
+Figure 3: Performance of Grad-Sim and Grad-Sim-Dr-Unsafe with varying target set size.
+Figure 3 illustrates the impact of target set variety on unsafe training data detection.
+For each target size, sampling is repeated 50 times, and both the mean and standard deviation are reported.
+We make two key observations.
+First, DRA achieves strong performance even with very limited target data: as few as 3 samples already enable it to outperform Wildguard.
+This demonstrates the practical effectiveness of our method in realistic low-resource scenarios.
+Second, DRA amplifies the benefits of increasing target set size.
+While raw gradient similarity (GradSim) struggles to fully leverage larger target sets due to noise in the gradient space, DRA denoises the representation, allowing the added diversity of unsafe examples to lead to improved detection performance.
+Appendix D $d^{\prime}$ as a Proxy for Downstream Metrics
+In this section we provide emphirical evidence that the $d^{\prime}$ we introduce in the main text match widely used detection metrics like AUPRC.
+Figure 4: Relationship between AUPRC and $d^{\prime}$, derived from data representation similarities
+Figure 4 illustrates the relationship between the AUPRC of the score computed from a single dimension and the corresponding $d^{\prime}$.
+The results show that AUPRC increases monotonically with $d^{\prime}$, and that $d^{\prime}$ serves as a reliable proxy for downstream metrics.
+This observation motivates our objective of maximizing $d^{\prime}$ via DRA.
+Appendix E Ablation Studies
+Table 7 reports ablations of our method.
+Applying either denoising (whitened PCA) or dimension selection based on $d^{\prime}$ obtained from the original space yields substantial gains over raw gradient similarity.
+Their combination achieves the best results across models and datasets showing that denoising and selection are complementary.
+Table 7: Ablation Studies for DRA across jailbreaking setup
+Model Method ToxicChat (%) XSTest-Response (%) $\boldsymbol{\Delta}$ Avg
+Llama-3-8B-Instruct GradSim (base) 36.4 1.7
+GradSim (Only Denoising) 60.1 21.7 +21.9+21.9
+GradSim (Only Dimension Selection) 50.3 65.5 +38.9+38.9
+GradSim-DRA 83.2 81.5 +63.3+63.3
+Gemma-2-9B-it GradSim (base) 41.1 50.2
+GradSim (Only Denoising) 61.1 64.4 +17.1+17.1
+GradSim (Only Dimension Selection) 59.2 62.3 +15.1+15.1
+GradSim-DRA 82.3 95.5 +43.3+43.3
+Appendix F Experimental Details
+For experiments involving model training, we train for 4 epochs with a warm up ratio of 0.1 and used learning rate of 1e-4 for Llama-3-8B-Instruct and 1e-5 for Gemma-2-9b-it respectively.
 The batch size is set to be 1 with no gradient accumulation.